### v2

# A Generalization of von Neumann’s Reduction from the Assignment Problem to Zero-Sum Games

Abstract
The equivalence between von Neumann’s Minimax Theorem for zero-sum games and the LP Duality Theorem connects cornerstone problems of the two fields of game theory and optimization, respectively, and has been the subject of intense scrutiny for seven decades. Yet, as observed in this paper, the proof of the difficult direction of this equivalence is unsatisfactory: It does not assign distinct roles to the two players of the game, as is natural from the definition of a zero-sum game.
In retrospect, a partial resolution to this predicament was provided in another brilliant paper of von Neumann (1953), which reduced the assignment problem to zero-sum games. However, the underlying LP is highly specialized — all entries of its objective function vector are strictly positive, the constraint vector is all ones, and the constraint matrix is 0/1.
We generalize von Neumann’s result along two directions, each allowing negative entries in certain parts of the LP. Our reductions make explicit the roles of the two players of the reduced game, namely their maximin strategies are to play optimal solutions to the primal and dual LPs. Furthermore, unlike previous reductions, the value of the reduced game reveals the value of the given LP. Our generalizations encompass several basic economic scenarios.
1 Introduction
The Minimax Theorem for zero-sum games (von Neumann, 1928), which is regarded as one of von Neumann’s spectacular contributions, not only started off [^1] Von Neumann is famously quoted as saying “As far as I can see, there could be no theory of games […] without that theorem […]. I thought there was nothing worth publishing until the Minimax Theorem was proved” (Casti, 1996, p. 19). game theory (von Neumann and Morgenstern, 1947) but also had far-reaching consequences in other areas. In particular, it is equivalent to the Linear Programming (LP) Duality Theorem and it led to Yao’s Lemma (Yao, 1977), a key tool for proving a lower bound on the worst-case performance of a randomized algorithm. The equivalence mentioned above is particularly significant since it connects cornerstone problems of the two fields of game theory and optimization. Our paper provides a fresh insight into this equivalence.
The two directions of this equivalence are qualitatively different: It is straightforward to reduce a zero-sum game to a primal-dual pair of LPs whose optimal solutions yield maximin strategies of the game (Dantzig, 1951). However, the other direction is much harder, i.e., the task of solving a linear program — or deciding that the primal or the dual program is infeasible — by finding maximin strategies of a corresponding zero-sum game. This direction follows from a series of papers published over a course of more than seven decades, starting with the seminal paper by Dantzig (1951). [^2] Interestingly, Dantzig (1982) describes a memorable meeting of his with von Neumann in which the latter outlined such a correspondence. Dantzig (1982) also attributes the LP Duality Theorem to von Neumann, again as a consequence of the same meeting, even though the first written proof is due to Tucker and his students Kuhn and Gale. However, the current reduction has an unsatisfactory aspect: it does not assign distinct roles to the two players of the reduced game, as is natural from the definition of a zero-sum game, see details below.
In retrospect, a partial resolution to this quandary was provided in another brilliant paper of von Neumann (1953), which reduced the assignment problem to zero-sum games; he called the latter the hide-and-seek game. Its drawback is that the underlying LP is highly specialized, see below. In this paper, we generalize von Neumann’s reduction along two directions, to two fairly large classes of LPs. Our generalizations capture several basic economic scenarios, beyond the assignment problem, and our reduction has the following properties:
  1. 1.
    It uses a scaled version of the constraint matrix as the game matrix, see details in the discussion following Theorem 4.3.
  2. 2.
    The row and column players obtain their maximin strategies by separately finding optimal solutions to the primal and dual LPs, respectively, thereby making explicit the roles of the two players.
  3. 3.
    If the LP has an optimal solution then the value of the reduced game is the reciprocal of the value of the LP.
We now examine in detail the current proofs of the two directions of the equivalence. For the easy direction, interpret the game matrix as the constraint matrix of a maximization LP and solve it optimally to obtain a maximin strategy for the row player (see, e.g., von Stengel, 2023, Section 2). Next optimally solve the dual LP to obtain a maximin strategy for the column player. By strong duality, the two LPs have the same objective function value, hence showing that the payoffs of the two players from these strategies must be equal. Observe that in general, the two players have distinct roles: their maximin strategies are solutions to different LPs and are vectors of different dimensions.
For the difficult direction, Dantzig (1951) reduced the given primal-dual pair of LPs to a symmetric zero-sum game; such a game has symmetric maximin strategies and the value of the game is zero. In this game, each player simultaneously mimics both primal and dual LPs, i.e., a strategy of the row player (or the column player) can be separated into two parts, corresponding to solutions of the primal and dual LPs. Even though this result stood for over half a century, it is in fact incomplete as observed by Adler (2013): the constructed zero-sum game can have solutions of a particular form which cannot be interpreted as optimal primal-dual pairs or certificates of infeasibility of the primal or the dual LP. Later — by adding an additional row to the zero-sum game — this reduction was extended to work in full generality, first by Adler (2013) and later, adding a certificate of infeasibility, by von Stengel (2023) and Brooks and Reny (2023); these are all strongly polynomial (in fact log-space), many-one reductions, see Section 3 for details. However, these extensions are valid only for LPs with rational numbers, hence giving the sense that the reduction is still incomplete, since the easy direction is valid for zero-sum games with real numbers (see Section 7.2 for a detailed discussion).
Next, we describe the partial resolution by von Neumann (1953). Let $(G,\mu)$ be an instance of the assignment problem on a complete bipartite graph $G$ with $n$ vertices on each side and strictly positive weights $\mu_{ij}$ on all edges $(i,j)$. The transformed game matrix $M$ is $2n\times n^{2}$, where the $2n$ rows correspond to the $2n$ vertices of $G$ and the $n^{2}$ columns correspond to the $n^{2}$ edges of $G$. Let $l$ represent edge $(i,j)$ of $G$. Then the $(k,l)^{th}$ entry of $M$ is defined as follows:
[…]
Observe that each column has exactly two nonzero entries.
The assignment problem asks for a maximum weight matching in bipartite graph $G$ under edge weights $\mu_{ij}$. In the hide-and-seek game, the row player picks one vertex and the column player picks one edge, say $e=(i,j)$, of $G$. If the vertex is one of the endpoints of the edge, then the column player pays $\frac{1}{\mu_{ij}}$ to the row player, and otherwise, no payment is made. The row player attempts to maximize his payoff and the column player minimizes the amount she has to pay. The reason for the name “hide-and-seek” should now be clear: the column player hides in an edge and the row player attempts to guess one of its endpoints. As usual, mixed strategies will be probability distributions over pure strategies.
Consider the primal-dual pair of LPs for the assignment problem in $G$: the primal LP finds a maximum weight fractional perfect matching in $G$ and the dual LP finds a minimum generalized vertex cover. The most remarkable step in the construction given above is von Neumann’s choice of the payoff matrix. Indeed, up to scaling, this matrix is exactly the constraint matrix of the LP. He shows that for this payoff matrix, the optimal solutions to the primal and dual LPs yield maximin strategies for the column and row players, respectively. But why is the payment the reciprocal of the edge weight? Our reduction generalizes the idea of using the scaled constraint matrix as the game matrix, and from this larger perspective, we are able to dispel the mystery behind von Neumann’s reduction, see Section 5.1.
A drawback of von Neumann’s reduction is that his LPs are highly specialized and have almost a “combinatorial” feel to them: His primal LP is $\max c^{\intercal}x\ \ \text{s.t. }Ax\leq\mathbf{1},\ x\geq 0$, with all entries in $c$ being strictly positive and the constraint matrix, $A$, being the incidence matrix of the underlying complete bipartite graph, i.e., all components of this matrix are zeros or ones. That still leaves a nagging doubt: can a separation of the roles of the two players be achieved if the LPs are more general, e.g., for a more general constraint vector $b$ or for data with negative entries. In this paper, we generalize von Neumann’s hide-and-seek game along two directions:
  1. 1.
    If $A$ is arbitrary, but $b$ and $c$ are strictly positive.
  2. 2.
    If $b$ and $c$ are arbitrary, but $A$ is nonnegative.
A wide range of applications are covered by these two extensions. The allocation of scarce resources, including classic economic scenarios such as the diet problem (via its dual LP) or farm planning, can be modelled with LPs which typically have all components of $A$, $b$, and $c$ positive (Chvátal, 1983, Chapter 11), as required by the first extension. Additional applications are covered by the second extension, e.g., the transportation problem which is equivalent to $b$-matching, a direct generalization of the assignment problem (Schrijver, 2003, Chapter 21).
In fact, the setting of LPs with positive $b$ and $c$ corresponds to production planning scenarios and gives rise to a natural game-theoretic interpretation extending von Neumann’s hide-and-seek game, as we discuss in Section 5.2. We call the resulting game the production planning game.
We leave the open problem of giving a complete generalization. We believe this will require a substantially new approach, since we seem to have extended von Neumann’s idea of rescaling as far as was possible. Our results include:
  1. 1.
    We cover larger classes of linear programs, beyond the assignment problem, that capture many important applications.
  2. 2.
    Our generalization even covers linear programs whose primal is unbounded (and hence dual is infeasible).
  3. 3.
    Our reduction carves out the exact reason why the hide-and-seek game uses the reciprocals of the assignment values for the payments.
  4. 4.
    Our generalization is valid for linear programs in real numbers.
Finally, since the reduced game always has a well-defined value, achieved by maximin strategies, it offers a new perspective on LP pairs, one of which is unbounded. The game-theoretic world allows us to distinguish such LPs; this can be interpreted as a degree of (in)feasibility, see Remark 4.4.
We end our paper by discussing evidence that von Neumann possessed an implicit understanding of the notion of polynomial time.
Related Work
Some previous work has proposed other generalizations of von Neumann’s hide-and-seek game. However, in contrast to our paper, none of these generalizations provides a reduction between linear programs and zero-sum games.
First, Marchi and Tarazaga (1981) generalized the hide-and-seek game to higher dimensions. Recall that the original interpretation of von Neumann’s hide-and-seek game is that the first player hides in one cell of a matrix, and the other player tries to find the first player by checking a row or a column of the matrix. Marchi and Tarazaga (1981) study a game in which the first player hides in a higher-dimensional tensor.
Second, Galichon and Jacquet (2024) study a different generalization of von Neumann’s result. In contrast to Marchi and Tarazaga (1981), their reduction is from a version of the assignment problem with a more general, nonlinear payoff structure. They reduce to a two-player game generalizing the hide-and-seek game that is not zero-sum anymore.
2 Preliminaries
We start with some conventions. Whenever we write an inequality between vectors, we assume that this inequality holds componentwise. Moreover, for a vector $v$ and a matrix $M$, we denote by $v^{\intercal}$ and $M^{\intercal}$ its transposed vector and matrix, respectively. Finally, we denote by $\mathbf{1}$ and $\mathbf{0}$ the vectors or matrices consisting of $1$- and 0-entries only; their dimension will be clear from the context.
We now introduce linear programs and zero-sum games, the main objects of our study. For text book introductions to these topics, we recommend the books by Schrijver (1998) and Nisan et al. (2007). We consider linear programs of the following form. Assume that we are given $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$. We denote their components by $a_{ij}$, $b_{i}$, and $c_{j}$, respectively. These objects specify the linear program $(P)$ together with its dual $(D)$.
$(P)$ $(D)$
$\text{max}\qquad c^{\intercal}x$ $\text{min}\qquad b^{\intercal}y$
$\text{s.t. }\qquad Ax$ $\leq b$ $\text{s.t. }\qquad A^{\intercal}y$ $\geq c$
$x$ $\geq 0$ $y$ $\geq 0$
We write $(P,D)$ to refer to the pair of linear programs $(P)$ and $(D)$. A vector $x\in\mathbb{R}^{n}$ is called (primal) feasible if it satisfies $Ax\leq b$ and $x\geq 0$, i.e., if it satisfies the constraints of $(P)$. Similarly, a vector $y\in\mathbb{R}^{m}$ is called (dual) feasible if it satisfies $A^{\intercal}y\geq c$ and $y\geq 0$, i.e., if it satisfies the constraints of $(D)$. The weak duality theorem states that if $x$ and $y$ are primal feasible and dual feasible, respectively, then $c^{\intercal}x\leq b^{\intercal}y$. This means that whenever we find feasible $x$ and $y$ such that $c^{\intercal}x=b^{\intercal}y$, then $x$ is a feasible vector maximizing $c^{\intercal}x$ and $y$ is a feasible vector minimizing $b^{\intercal}y$. In this case, we say that $(x,y)$ an optimal primal-dual pair. The strong duality theorem states that whenever $(P)$ and $(D)$ admit feasible points, then they admit an optimal primal-dual pair, and if so, $c^{\intercal}x=b^{\intercal}y$ is called the value of the LP. The following characterization of possible outcomes of a pair of primal and dual programs $(P,D)$ is well-known.
Proposition 2.1.
For any pair of linear programs $(P,D)$, exactly one of the following outcomes is true:
  1. 1.
    There exist $x^{*},y^{*}$ such that
    $Ax^{*}\leq b,\;x^{*}\geq\mathbf{0};\;A^{\intercal}y^{*}\geq c,\;y^{*}\geq\mathbf{0},\;c^{\intercal}x^{*}=b^{\intercal}y^{*}.$
    In this case $x^{*}$ and $y^{*}$ are optimal solutions to $(P)$ and $(D)$ respectively.
    There exists $z^{*}$ such that
    $A^{\intercal}z^{*}\geq\mathbf{0},\;z^{*}\geq\mathbf{0},\;b^{\intercal}z^{*}<0.$
    In this case $(P)$ is infeasible. We call $z^{*}$ an infeasibility certificate for $(P)$.
  2. 2.
    There exists $w^{*}$ such that
    $Aw^{*}\leq\mathbf{0},\;w^{*}\geq\mathbf{0},\;c^{\intercal}w^{*}>0.$
    In this case $(P)$ is unbounded if feasible. We call $w^{*}$ an unboundedness certificate for $(P)$.
Given $A\in\mathbb{R}^{m\times n}$, $c\in\mathbb{R}^{n}$, and $b\in\mathbb{R}^{m}$, the problem of solving the linear program $(P)$ is to provide one of the three outputs stated above.
Remark 2.2.
We make a few important remarks.
In order to certify $(P)$ to be unbounded, we need to show, in addition to the unboundedness certificate, a feasible solution to $(P)$ (i.e., a vector $x$ such that $Ax\leq b$, $x\geq\mathbf{0})$.
If $z^{*}$ is an infeasibility certificate for $(P)$, then $z^{*}$ is an unboundedness certificate for $(D)$.
If $w^{*}$ is an unboundedness certificate for $(P)$, then $w^{*}$ is an infeasibility certificate for $(D)$.
Cases (2) and (3) in Proposition 2.1 establish that the linear programs have no optimal solutions by certifying that either $(P)$ or $(D)$ have no feasible solution.
Our second important concept is zero-sum games. A zero-sum game is specified by a matrix $M=(m_{ij})\in\mathbb{R}^{m\times n}$. There are two players, a row player and a column player; for ease of notation, we will assume that the first is male and the second is female. The row player’s action is to select a row $i$ with $1\leq i\leq m$ and the column player’s action is to select a column $j$ with $1\leq j\leq n$. Selecting row $i$ and column $j$ leads to a payment of the associated matrix entry, $m_{ij}$, from the column player to the row player. Hence, whereas the row player’s goal is to maximize the value of the matrix entry chosen, the column player’s goal is to minimize it.
The players are allowed to randomize their actions. A mixed strategy is a probability distribution over actions. The row and column players select mixed strategies over rows and columns, respectively. These can be specified by vectors $p\in\mathbb{R}^{m}$ and $q\in\mathbb{R}^{n}$ that are nonnegative and satisfy $p^{\intercal}\mathbf{1}=q^{\intercal}\mathbf{1}=1$. The entries of $p$ and $q$ are the probabilities with which the respective rows and columns are played. If the row player plays $p$ and the column player plays $q$, then the row player receives the expected payment of $p^{\intercal}Mq$ from the column player.
As a solution concept for zero-sum games, we consider maximin strategies. These are strategies that give a best possible guarantee to a player, even if the opponent plays the worst possible counter-strategy. More precisely, a maximin strategy for the row player (who tries to maximize his payoff) is a strategy $p^{*}$ that satisfies
$p^{*}\in\arg\max_{p}\min_{q}p^{\intercal}Mq\text{.}$
Similarly, a minimax strategy for the column player is a strategy $q^{*}$ that satisfies
$q^{*}\in\arg\min_{q}\max_{p}p^{\intercal}Mq$
If we replace the matrix $M$ by its negative, we see that this is the same as a maximin strategy in the game where the column player tries to maximize the negative payoff she receives. Therefore, we simplify terminology and say that we compute maximin strategies for both players when we actually mean computing a maximin strategy for the row player and a minimax strategy for the column player. We call $(p^{*},q^{*})$ a pair of maximin strategies.
The unique payments determined by these strategies, i.e., $u^{*}=\max_{p}\min_{q}p^{\intercal}Mq$ and $v^{*}=\min_{q}\max_{p}p^{\intercal}Mq$ are called the maximin value and minimax value of the game, respectively. The famous Minimax Theorem by von Neumann (1928) states that
$u^{*}=\max_{p}\min_{q}p^{\intercal}Mq=\min_{q}\max_{p}p^{\intercal}Mq=v^{*}\text{,}$
i.e., the maximin value of the game coincides with its minimax value. Therefore, this quantity is simply called the value of the game. The goal of solving a zero-sum game is to compute a pair of maximin strategies. Note that such strategies are mutual best responses, and therefore constitute a Nash equilibrium.
It is well known that maximin strategies as well as the value of the game can be computed by the following linear programs; this is the easy direction of the correspondence between linear programs and zero-sum games (see, e.g., von Stengel, 2023, Equations (7) and (8)):
$($ $\mbox{{Row}})$ $~($ $\mbox{{Col}})$
max $u$ min $v$
s.t. $M^{\intercal}p\geq\mathbf{1}u$ s.t. $Mq\leq\mathbf{1}v$
$\mathbf{1}^{\intercal}p=1,\;p\geq\mathbf{0}$ $\mathbf{1}^{\intercal}q=1,\;q\geq\mathbf{0}$
Note that $(\mbox{{Row}},\mbox{{Col}})$ are a primal-dual pair. The optimal value of the linear program $(\mbox{{Row}})$ is the maximin value corresponding to the row player. An optimal solution $(u^{*},p^{*})$ of $(\mbox{{Row}})$ consists of the maximin value together with a maximin strategy. In fact, the constraints of $(\mbox{{Row}})$ imply that $p$ is a probability distribution and that $p^{\intercal}M\geq\mathbf{1}^{\intercal}u$. Therefore, ${p^{*}}^{\intercal}Mq\geq u^{*}$ for all mixed strategies $q$ of the column player. Similarly, $(\mbox{{Col}})$ determines the minimax value and a minimax strategy of the column player.
3 Reducing Linear Programs to Zero-Sum Games
In this section, we lay the groundwork for performing the difficult direction of the correspondence, i.e., reductions from LPs to zero-sum games. We start by reviewing the ideas behind the original reduction by Dantzig (1951).
In his seminal paper already discussed in the introduction, Dantzig (1951) introduced a reduction from LPs to zero-sum games by using a skew-symmetric game matrix. More precisely, given the primal LP $(P)$, he considers the zero-sum game with payoff matrix [^3] Dantzig (1951) attributes the idea for this matrix to Tucker.
$\begin{pmatrix}\vskip 6.0pt plus 2.0pt minus 2.0pt\mathbf{0}&A&-b\\ -A^{\intercal}&\mathbf{0}&c\\ b^{\intercal}&-c^{\intercal}&0\\ \end{pmatrix}\text{.}$ (1)
As this matrix is skew-symmetric (so it is a symmetric game), the value of the game is 0. Denoting a maximin strategy as $(p^{*},q^{*},t^{*})$, the key result of the paper by Dantzig (1951) states that
  1. 1.
[…]
  2. 2.
    If $b^{\intercal}p^{*}-c^{\intercal}q^{*}<0$, then either $z^{*}=p^{*}$ satisfies (2) or $w^{*}=q^{*}$ satisfies (3) in Proposition 2.1.
Note that since $(p^{*},q^{*},t^{*})$ is a maximin strategy for a game of value 0, it holds that $t^{*}(b^{\intercal}p^{*}-c^{\intercal}q^{*})=0$. The goal of solving $(P)$ is achieved, except for the case $t^{*}=b^{\intercal}p^{*}-c^{\intercal}q^{*}=0$. Notably, this “hole” in the reduction was clearly recognized by Dantzig (1951).
Incredibly, Dantzig’s reduction was nonetheless accepted and ascribed as complete for fifty years. Only then, a series of papers have worked towards a complete reduction. First, Adler (2013) modified Dantzig’s reduction and proposed a many-one reduction for the case that $A$, $b$, and $c$ are rational numbers. In the sequel, both Brooks and Reny (2023) and von Stengel (2023) noted that the output of Adler’s reduction was either a pair $(x^{*},y^{*})$ satisfying (1) in Proposition 2.1 or an assurance that there exists either $z^{*}$ or $w^{*}$ (possibly both) satisfying (2) or (3) in Proposition 2.1. Both papers introduce modifications adding a specific certificate of infeasibility that allows for solving $(P)$ as stated in Proposition 2.1.
The key idea of these papers is to modify the reduced game by Dantzig: Adler (2013) and von Stengel (2023) add an additional row while Brooks and Reny (2023) transform the game with elementary column operations. These reductions have in common the fact that the constraint matrix or vectors in the reduced games need large numbers, exponential in $m$ or nn. [^4] As an example, in the reduction by Adler (2013), the additional entry of the constraint vector is estimated as a number that is at least $m(\sqrt{m}\theta)^{m}+1$ where $\theta$ is the absolute value of the largest number contained in $A$ and $b$. A second way to solve linear programs explored by Adler (2013) is to use the original reduction by Dantzig and sequentially eliminate variables until the solution is in the right form. While this avoids large numbers, and is applicable to real numbers, it requires solving more than one zero-sum game. In other words, it is a Turing reduction rather than a many-one reduction. [^5] We refer to Section 7 for an in-depth discussion of the computational properties of our reduction.
4 Extending Von Neumann’s Reduction
In this section, we provide our extension of von Neumann’s reduction. First, we consider the case that the vectors $b>\mathbf{0}$ and $c>\mathbf{0}$, i.e., both are strictly positive. Next we will show how this enables us to tackle the case that the constraint matrix is nonnegative, i.e., $A\geq\mathbf{0}$.
4.1 Positive Objective and Constraint Vectors
The key technical step of our reduction is a rescaling of the linear programs $(P)$ and $(D)$; this is where we require $b>\mathbf{0}$ and $c>\mathbf{0}$, i.e., both vectors need to be strictly positive. We consider the diagonal matrices $B\in\mathbb{R}^{m\times m}$ and $C\in\mathbb{R}^{n\times n}$ defined by
[…]
We define the matrix $M=BAC\in\mathbb{R}^{m\times n}$ (its name already suggests that this will be the payoff matrix of our reduced zero-sum game) and a related pair of linear programs as follows.
$(P^{\prime})$ $(D^{\prime})$
$\text{max}\qquad\mathbf{1}^{\intercal}\xi$ $\text{min}\qquad\mathbf{1}^{\intercal}\eta$
$\text{s.t. }\qquad M\xi$ $\leq\mathbf{1}$ $\text{s.t. }\qquad M^{\intercal}\eta$ $\geq\mathbf{1}$
$\xi$ $\geq\mathbf{0}$ $\eta$ $\geq\mathbf{0}$
The relationship with the original linear programs is captured as follows.
Proposition 4.1.
Assume that $b>\mathbf{0}$ and $c>\mathbf{0}$. The following statements are true:
  1. 1.
    1. (a)
      $\xi$ is feasible for $(P^{\prime})$ if and only if $C\xi$ is feasible for $(P)$,
    2. (b)
      $\eta$ is feasible for $(D^{\prime})$ if and only if $B\eta$ is feasible for $(D)$.
  2. 2.
    $(\xi^{*},\eta^{*})$ is an optimal primal-dual pair for $(P^{\prime})$ and $(D^{\prime})$ if and only if $(C\xi^{*},B\eta^{*})$ is an optimal primal-dual pair for $(P)$ and $(D)$.
  3. 3.
    $\bar{\xi}$ is an unboundedness certificate for $(P^{\prime})$ if and only if $C\bar{\xi}$ is an unboundedness certificate for $(P)$.
Proof.
We prove the statements one by one.
  1. 1.
    Clearly, $\xi\geq 0$ if and only if $C\xi\geq 0$. Moreover, $M\xi\leq\mathbf{1}$ if and only if $A(C\xi)=B^{-1}M\xi\leq B^{-1}\mathbf{1}=b$. This implies (a). The proof of (b) is analogous.
  2. 2.
    From ($a$) and ($b$) we have that $\xi,\eta$ are feasible for $(P^{\prime}),(D^{\prime})$ if and only if $C\xi,B\eta$ are feasible for $(P),(D)$, respectively. In addition, $c^{\intercal}(C\xi)=\mathbf{1}^{\intercal}\xi$ and $b^{\intercal}(B\eta)=\mathbf{1}^{\intercal}\eta$. Applying proposition 2.1, completes the proof.
  3. 3.
    This follows from (1) and the observation that, because $c>\mathbf{0}$, $\mathbf{1}^{\intercal}\xi>0$ if and only if $\mathbf{1}^{\intercal}C\xi>0$. Hence, $\bar{\xi}\in\{\xi\;|\;M\xi\leq\mathbf{0},\;\xi\geq\mathbf{0},\;\mathbf{1}^{\intercal}\xi>0\}$ if and only if $C\bar{\xi}\in\{x\;|\;Ax\leq\mathbf{0},\;x\geq\mathbf{0},\;\mathbf{1}^{\intercal}x>0\}$.∎
Remark 4.2.
Proposition 4.1 covers cases (1) and (2) of Proposition 2.1, i.e., there exists an optimal primal-dual pair of solutions for $(P)$ or that $(P)$ is unbounded. Case (LABEL:it:inf), when $(P)$ and $(P^{\prime})$ are infeasible, never occurs under the assumption that $b>\mathbf{0}$, since in this case, $x=\mathbf{0}$ is a primal feasible solution.
Let $(G)$ be the zero-sum game with payoff matrix $M$ where the row-player wants to maximize his payoff and the column player wants to minimize her payoff. Recall that the maximin and minimax values of this game correspond to the optimal values of the linear programs introduced in Section 2:
$($ $\mbox{{Row}})$ $~($ $\mbox{{Col}})$
max $u$ min $v$
s.t. $M^{\intercal}p\geq\mathbf{1}u$ s.t. $Mq\leq\mathbf{1}v$
$\mathbf{1}^{\intercal}p=1,\;p\geq 0$ $\mathbf{1}^{\intercal}q=1,\;q\geq 0$
We have the following theorem.
Theorem 4.3.
Assume that $b>\mathbf{0}$ and $c>\mathbf{0}$.
  1. 1.
    Let $(p^{*},q^{*})$ be a pair of maximin strategies of the zero-sum game $(G)$ and $v^{*}(=u^{*})$ its value.
    1. (a)
      If $v^{*}>0$, then $\left(\frac{1}{v^{*}}q^{*},\frac{1}{v^{*}}p^{*}\right)$ is an optimal primal-dual pair for $(P^{\prime})$ and $(D^{\prime})$.
    2. (b)
      If $v^{*}\leq 0$, then $q^{*}$ is an unboundedness certificate for $(P^{\prime})$.
  2. 2.
    Let $v^{*}$ be the value of the game $(G)$.
    1. (a)
      If $(\xi^{*},\eta^{*})$ is an optimal primal-dual pair for $(P^{\prime})$ and $(D^{\prime})$, then ${\mathbf{1}^{\intercal}\xi^{*}}\left(={\mathbf{1}^{\intercal}\eta^{*}}\right)>0$, $\left(\frac{1}{\mathbf{1}^{\intercal}\eta^{*}}\eta^{*},\frac{1}{\mathbf{1}^{\intercal}\xi^{*}}\xi^{*}\right)$ is a pair of maximin strategies of the zero-sum game $(G)$, and $v^{*}=\frac{1}{\mathbf{1}^{\intercal}\xi^{*}}$.
    2. (b)
      If $w^{*}$ is an unboundedness certificate of $(P^{\prime})$, then $v^{*}\leq 0$.
Proof.
We prove the statements one by one.
  1. 1-(a)
    The feasibility of $\left(\frac{1}{v^{*}}q^{*},\frac{1}{v^{*}}p^{*}\right)$ for $(P^{\prime})$ and $(D^{\prime})$ is readily verified by the feasibility of $(p^{*},q^{*})$ for $(\mbox{{Col}})$ and $(\mbox{{Row}})$. In addition, $\mathbf{1}^{\intercal}p^{*}=\mathbf{1}^{\intercal}q^{*}=1$ implies $\mathbf{1}^{\intercal}\frac{1}{v^{*}}q^{*}=\mathbf{1}^{\intercal}\frac{1}{v^{*}}p^{*}$ which completes the proof.
  2. 1-(b)
    Suppose that $v^{*}\leq 0$. Then, $Mq^{*}\leq\mathbf{1}v^{*}\leq\mathbf{0}$, $\mathbf{1}^{\intercal}q^{*}=1>0$, and $q^{*}\geq\mathbf{0}$. Hence, $q^{*}$ is an unboundedness certificate for $(P^{\prime})$.
  3. 2-(a)
    First, note that the optimal objective value of $(P^{\prime})$ is always strictly larger than 0, and, therefore, $\mathbf{1}^{\intercal}\xi^{*}>0$. Define $\hat{v}:=\frac{1}{\mathbf{1}^{\intercal}\xi^{*}}>0$. We have that $\mathbf{1}^{\intercal}(\hat{v}\eta^{*})=\mathbf{1}^{\intercal}(\hat{v}\xi^{*})=1$, $\hat{v}\eta^{*}\geq 0$, and $\hat{v}\xi^{*}\geq 0$, and, therefore, $\hat{v}\eta^{*}$ and $\hat{v}\xi^{*}$ are valid strategies.
    Next, we show that $\hat{v}\eta^{*}$ and $\hat{v}\xi^{*}$ are mutual best responses. Consider any strategy $q$ of the column player. By dual feasibility, it holds that
    $(\hat{v}\eta^{*})^{\intercal}Mq\geq(\hat{v}\mathbf{1})^{\intercal}q=\hat{v}\text{.}$ (2)
    There, $(\hat{v}\mathbf{1})^{\intercal}q=\hat{v}$ holds because $q$ is a probability distribution. Moreover, by primal feasibility,
    $(\hat{v}\eta^{*})^{\intercal}M(\hat{v}\xi^{*})\leq(\hat{v}\eta^{*})^{\intercal}\hat{v}\mathbf{1}=\hat{v}\hat{v}{\eta^{*}}^{\intercal}\mathbf{1}=\hat{v}\hat{v}\frac{1}{\hat{v}}=\hat{v}\text{.}$ (3)
    Hence, $\hat{v}\eta^{*}$ is a best response for the row player. The computation for the column player is analogous. Together, $(\hat{v}\eta^{*},\;\hat{v}\xi^{*})$ is a Nash equilibrium and therefore a pair of maximin strategies for $(G)$.
    Finally, combining Equation 2 for $q=\hat{v}\xi^{*}$ and Equation 3, we know that $(\hat{v}\eta^{*})^{\intercal}M(\hat{v}\xi^{*})=\hat{v}$. Hence, $v^{*}=\hat{v}$ is the value of the game $(G)$.
  4. 2-(b)
    Assume that $w^{*}$ is an unboundedness certificate for $(P^{\prime})$, i.e., $Mw^{*}\leq\mathbf{0}$, $w^{*}\geq\mathbf{0}$, and $\mathbf{1}^{\intercal}w^{*}>0$. Let $\alpha=\frac{1}{\mathbf{1}^{\intercal}w^{*}}>0$ and consider $q=\alpha w^{*}$ and $v=0$. We claim that $(q,v)$ is feasible for $(\mbox{{Col}})$.
    Indeed, since $w^{*}\geq 0$, it holds that $\alpha w^{*}\geq 0$. In addition, $\mathbf{1}^{\intercal}q=\alpha\mathbf{1}^{\intercal}w^{*}=1$. Moreover, $Mq=\alpha Mw^{*}\leq\alpha\mathbf{0}=\mathbf{1}v$.
    Hence, $(q,v)$ is feasible for $(\mbox{{Col}})$ and therefore $v^{*}\leq 0$. ∎
An appealing aspect of our reduction is that it uses a scaled version of the constraint matrix, $A$, as the game matrix. The scaling is done as follows: the columns of $A$ are scaled by the reciprocals of the coefficients in the objective function, $c$, and the rows of $A$ are scaled by the reciprocals of the coefficients in the right hand side, $b$. In addition to being natural, this approach strengthens the connection between the original linear program and the reduced game beyond what was achieved by Dantzig and his successors.
In particular, the games by Dantzig and Adler are symmetric and have value zero, i.e., the value of the game has no information. [^6] While we regard the nonsymmetry of our reduced games as beneficial for interpreting the game, we acknowledge that reductions producing symmetric games have other, computational advantages. For instance, the symmetrization techniques by Gale et al. (1950) imply that algorithms for solving symmetric games can be applied to nonsymmetric games without significant further computational effort. The value of the reduced games by von Stengel as well as Brooks and Reny are zero if and only if there exists an optimal solution to the original LP. In contrast, the value of our reduced game has valuable information: it allows us to decide whether the LP admits an optimal solution and, if this is the case, it is the reciprocal of its optimal value. We believe this is a potent fact which may find a suitable application.
We conclude this subsection with a remark about how to interpret the value of the reduced game as a degree of (in)feasibility.
Remark 4.4.
The primal-dual pair $(\mbox{{Row}},\mbox{{Col}})$ always has primal and dual feasible solutions, whereas $(D^{\prime})$ might be infeasible. Recall from Remark 2.2 that an infeasibility certificate for $(D^{\prime})$ is equivalent to an unboundedness certificate for $(P^{\prime})$, i.e., we ask whether there exists $w^{*}$ with $Mw^{*}\leq\mathbf{0}$, $w^{*}\leq 0$, and $\mathbf{1}^{\intercal}w^{*}>0$. By scaling $w^{*}$, the last condition can equivalently be written as $\mathbf{1}^{\intercal}w^{*}=1$. The LP $(\mbox{{Col}})$ only differs from these constraints by adding the minimum necessary slack to the constraints governed by $v$. Hence, the value of the reduced zero-sum game can be interpreted as a degree of feasibility, i.e., the minimum slack needed to guarantee feasibility of $(D^{\prime})$. Games with a nonpositive value indicate infeasibility of $(D^{\prime})$, and their exact value can then be seen as measuring a degree of infeasibility.
4.2 Nonnegative Constraint Matrix
Finally, we show that the correspondence described in Theorem 4.3 is applicable in a different setting as well. We demonstrate how to use the developed theory to deal with linear program pairs $(P,D)$ when $A\geq\mathbf{0}$ but $b$ and $c$ are not constrained to be positive.
The intuitive idea is to preprocess the linear program to satisfy the conditions that $b$ and $c$ are positive: If $b_{i}<0$, then the nonnegativity of $A$ immediately implies infeasibility. Otherwise, we want to delete all columns and rows that correspond to nonpositive entries of $b$ and $c$. If $b_{i}=0$, then $x_{j}=0$ whenever there exists $j\in[n]$ with $a_{ij}>0$. Hence, we can delete the $i$th row together with all such columns, and later on insert $x_{j}=0$ when reconstructing solutions. Similarly, $c_{j}\leq 0$ requires $x_{j}=0$ and we can omit such columns to transition to a smaller LP. We thus reach an LP of the original form, for which we can apply the correspondence of Section 4.1.
We now define sets $I_{0}$, $J_{0}$, and $J_{-}$ of indices of rows and columns to be deleted according to the above discussion.
$I_{0}$ $:=\{i\in[m]\;|\;b_{i}=0\},$
$J_{0}$ $:=\{j\in[n]\;|\;a_{ij}>0\text{ for some }i\in I_{0}\},\textnormal{ and }$ (4)
$J_{-}$ $:=\{j\in[n]\;|\;c_{j}\leq 0\}$
[…]
$\text{max}\qquad\hat{c}^{\intercal}x$ $\text{min}\qquad\hat{b}^{\intercal}y$
$\text{s.t. }\qquad\hat{A}x$ $\leq\hat{b}$ $\text{s.t. }\qquad\hat{A}^{\intercal}y$ $\geq\hat{c}$
$x$ $\geq 0$ $y$ $\geq 0$
[…]
We collect the key insights for the correctness of Algorithm 1 in the following theorem.
Input: Linear program pair specified by $A\geq\mathbf{0}$, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$
Output: Certificate according to Proposition 2.1
1
2if there exists $i\in[m]$ such that $b_{i}<0$ then
3   Define $z^{*}$ by $z^{*}_{k}=\begin{cases}1&k=i\\ 0&k\neq i\end{cases}$.
4
5   return Infeasibility certificate $z^{*}$
6Define $I_{0}$, $J_{0}$, and $J_{-}$ as in Equation 4.
7 if there exists $j\in[n]\setminus J_{-}$ such that $a_{ij}=0$ for all $i\in[m]$ then
8   Define $w^{*}$ by $w^{*}_{k}=\begin{cases}1&k=j\\ 0&k\neq j\end{cases}$.
9
10   return Unboundedness certificate $w^{*}$
11Define $\lambda:=\max\left\{\left.\frac{c_{j}}{a_{ij}}\;\right|\;j\in J_{0},i\in I_{0}\text{ with }a_{ij}>0\right\}$ if $J_{0}\neq\emptyset$ and $\lambda=0$ if $J_{0}=\emptyset$.
12 if $I_{0}=[m]$ or $J_{0}\cup J_{-}=[n]$ then
13    Define $(x^{*},y^{*})$ by $x^{*}=\mathbf{0}$ and $y^{*}_{i}:=\begin{cases}\lambda&i\in I_{0}\\ 0&i\notin I_{0}\end{cases}$.
14    return Optimal primal-dual pair $(x^{*},y^{*})$
[…]
17    Define $(x^{*},y^{*})$ by $x^{*}_{j}:=\begin{cases}0&j\in J_{0}\cup J_{-}\\ \hat{x}_{j}&j\notin J_{0}\cup J_{-}\end{cases}$ and $y^{*}_{i}:=\begin{cases}\lambda&i\in I_{0}\\ \hat{y}_{i}&i\notin I_{0}\end{cases}$.
18    return Optimal primal-dual pair $(x^{*},y^{*})$
Algorithm 1 Solving linear programs with $A\geq\mathbf{0}$
Theorem 4.5.
Assume that $A\geq\mathbf{0}$.
  1. 1.
    $(P)$ is feasible if and only if $b\geq\mathbf{0}$. If there exists $i\in[m]$ such that $b_{i}<0$, then Algorithm 1 returns an infeasibility certificate for $(P)$ in algorithm 1.
  2. 2.
    Suppose that $b\geq\mathbf{0}$.
    1. (a)
      $(P)$ is unbounded if and only if there exists $j\in[n]\setminus J_{-}$ such that $a_{ij}=0$ for all $i\in[m]$. In this case, Algorithm 1 returns an unboundedness certificate for $(P)$ in algorithm 1.
    2. (b)
      If $I_{0}=[m]$ or $J_{0}\cup J_{-}=[n]$, then Algorithm 1 returns an optimal primal-dual pair for $(P,D)$ in algorithm 1.
  3. 3.
[…]
Proof.
We prove the statements one by one.
1.
If $b\geq\mathbf{0}$, then $x=\mathbf{0}$ is a feasible solution of $(P)$. Conversely, if $x$ is a feasible solution of $(P)$, then $b\geq Ax\geq\mathbf{0}$, where the second inequality is true because $A\geq\mathbf{0}$ and $x\geq\mathbf{0}$.
Moreover, assume that there exists $i\in[m]$ such that $b_{i}<0$ and consider $z^{*}$ as defined in Algorithm 1. Then $A^{\intercal}z^{*}\geq\mathbf{0}$, $z^{*}\geq\mathbf{0}$, and $b^{\intercal}z^{*}<0$. Hence, $z^{*}$ is an infeasibility certificate for $(P)$.
2-(a)
Assume that there exists $j\in[n]\setminus J_{-}$ such that $a_{ij}=0$ for all $i\in[m]$. Consider $w^{*}$ as returned in algorithm 1. Clearly $w^{*}\geq\mathbf{0}$. In addition, since $j\notin J_{-}$, it holds that $c_{j}>0$ and therefore $c^{\intercal}w^{*}>0$. Finally, for every $i\in[m]$ it holds that every $(Aw^{*})_{i}=a_{ij}=0$. We used that only the $j$th entry of $w^{*}$ is nonzero and our assumption on the entries of $A$.
Conversely, assume that $(P)$ admits an unboundedness certificate $w^{*}$, i.e., $w^{*}\geq\mathbf{0}$, $Aw^{*}\leq\mathbf{0}$, and $c^{\intercal}w^{*}>0$. Since $c^{\intercal}w^{*}>0$ and $w^{*}\geq\mathbf{0}$, there exists $j^{*}\in[n]$ with $c_{j^{*}}>0$ and $w^{*}_{j^{*}}>0$. By definition $j^{*}\notin J_{-}$. Define $A_{j^{*}}$ as the $j^{*}$th column of $A$. Then, $\mathbf{0}\geq Aw^{*}\geq w^{*}_{j^{*}}A_{j^{*}}$. Since $w^{*}_{j^{*}}>0$, it follows that $A_{j^{*}}\leq\mathbf{0}$ and therefore $A_{j^{*}}=\mathbf{0}$ by the nonnegativity of $A$.
2-(b)
Consider the pair $(x^{*},y^{*})$ returned in algorithm 1. Clearly $x^{*}=\mathbf{0}$ is primal feasible. Moreover, $y^{*}\geq 0$. Now let $j\in[n]$. If $j\in J_{-}$, the $j$th constraint is satisfied. Otherwise, it holds that $j\in[n]\setminus J_{-}$ and we have already excluded the case where $a_{ij}=0$ for all $i\in[m]$. Hence, there exists $i\in[m]$ with $a_{ij}>0$. By definition of $\lambda$, it holds that $\lambda\geq\frac{c_{j}}{a_{ij}}$. Therefore, $(A^{\intercal}y^{*})_{j}\geq a_{ij}y^{*}_{i}=a_{ij}\lambda\geq a_{ij}\frac{c_{j}}{a_{ij}}=c_{j}$. Hence, $y^{*}$ is dual feasible. Finally, it holds that $b^{\intercal}y^{*}=\mathbf{0}=c^{\intercal}x^{*}$ where the first inequality holds by definition of $y^{*}$ (whose components are only nonzero for indices in $I_{0}$). Together, $(x^{*},y^{*})$ is an optimal primal-dual pair.
3.
We will first prove the existence of feasible solutions to $\hat{P}$ and $\hat{D}$. Clearly $\hat{x}=\mathbf{0}$ is feasible for (P^)($\hat${P}) because $\hat{b}>\mathbf{0}$.
Moreover, since $\hat{b}\geq 0$, we can apply part 2-(a) to conclude that (P^)($\hat${P}) is bounded if $\hat{A}$ does not contain 0-columns (which is a stronger condition than only requiring certain nonzero columns). Let $j\in[n]\setminus(J_{0}\cup J_{-})$. We claim that there exists $i\in[m]\setminus I_{0}$ with $a_{ij}>0$. Assume that this was not the case, i.e., $a_{ij}=0$ for all $i\in[m]\setminus I_{0}$. Note that since $j\notin J_{0}$, we also have that $a_{ij}=0$ for all $i\in I_{0}$. However, this would imply that we have identified a $j\in[n]\setminus J_{-}$ such that $a_{ij}=0$ for all $i\in[m]$, contradicting that Algorithm 1 has reached algorithm 1. Hence, (P^)($\hat${P}) is feasible and bounded, and, therefore, (D^)($\hat${D}) is feasible as well. By strong LP duality, we have the existence of an optimal primal-dual pair.
[…]
Moreover, for all indices $j\in J_{-}$, it holds that $(A^{\intercal}y^{*})_{j}\geq 0\geq c_{j}$ where we use the nonnegativity of $A$ and $y^{*}$ for the first inequality. Additionally, if $j\in J_{0}$, then there exists $i\in I_{0}$ with $a_{ij}>0$. By definition of $\lambda$, it holds that $\lambda\geq\frac{c_{j}}{a_{ij}}$. Hence, $(A^{\intercal}y^{*})_{j}\geq a_{ij}y^{*}_{i}=a_{ij}\lambda\geq a_{ij}\frac{c_{j}}{a_{ij}}=c_{j}$. It follows that $y^{*}$ is dual feasible.
Finally, it holds that
$c^{\intercal}x^{*}=\hat{c}^{\intercal}\hat{x}=\hat{b}^{\intercal}\hat{y}=b^{\intercal}y^{*}\text{.}$ (5)
[…]
Corollary 4.6.
Algorithm 1 solves linear programs with $A\geq\mathbf{0}$.
5 Application and Interpretation
In this section, we want to demonstrate the applicability of our reductions. We start by deepening our understanding of von Neumann’s hide-and-seek game in the context of our more general reduction. We then discuss playing the reduced game for any LP with positive $b$ and $c$. Finally, we provide an interpretation of the reduced game in relationship to its original LP, where we interpret the players’ strategies as “playing” the primal and dual LP.
5.1 Von Neumann’s Correspondence for the Assignment Problem
Originally, von Neumann had designed his reduction for the special case where the source LP corresponds to the (fractional) assignment problem. The assignment problem consists of $n$ workers and $n$ jobs and it is assumed that assigning worker $i$ to job $j$ creates a positive surplus $\mu_{ij}>0$. A fractional assignment of workers to jobs is a nonnegative matrix $(x_{ij})$, where
  • for all $1\leq i\leq n$, it holds that $\sum_{j=1}^{n}x_{ij}\leq 1$, and
  • for all $1\leq j\leq n$, it holds that $\sum_{i=1}^{n}x_{ij}\leq 1$.
The constraints ensure that each worker is assigned no more than the fractions equivalent to one job, and no job is assigned more than one worker. The goal is to find a fractional assignment that maximizes the expected surplus. [^7] Due to the famous algorithm by Birkhoff (1946), an optimal fractional assignment can be decomposed into perfect matchings of the same value. Hence, finding an optimal fractional assignment is sufficient to find an optimal integral assignment. In fact, the paper by von Neumann (1953) is more famous as a reference for the Birkhoff-von Neumann Decomposition Theorem of bistochastic matrices rather than for introducing the hide-and-seek game. This is equivalent to solving the linear program $(\mbox{{Asg}})$ whose dual is given by $(\overline{\mbox{{Asg}}})$.
$(\mbox{{Asg}})$
$\text{max}\qquad\sum_{i,j=1}^{n}\mu_{ij}x_{ij}$
$\text{s.t. }\qquad\sum_{j=1}^{n}x_{ij}$ $\leq 1\ \forall\,i\in[n]$
$\qquad\sum_{i=1}^{n}x_{ij}$ $\leq 1\ \forall\,j\in[n]$
$x_{ij}$ $\geq 0\ \forall\,i,j\in[n]$
$(\overline{\mbox{{Asg}}})$
$\text{min}\qquad\sum_{i=1}^{n}u_{i}+\sum_{j=1}^{n}v_{j}$
$\text{s.t. }\qquad u_{i}+v_{j}$ $\geq\mu_{ij}\ \forall\,i,j\in[n]$
$u_{i}$ $\geq 0\ \forall\,i\in[n]$
$v_{j}$ $\geq 0\ \forall\,j\in[n]$
(Asg) is a linear program of the form $(P)$ where the objective vector is strictly positive (containing the entries $\mu_{ij}$), the constraint matrix contains only zeros and ones, and the right hand side is the all-ones vector. Its dual contains variables $u_{i}$ and $v_{j}$ corresponding to workers and jobs of the assignment problem.
Now, matrix $B$ for $(\mbox{{Asg}})$ is the identity matrix, and matrix $C$ is the diagonal matrix of dimension $n^{2}\times n^{2}$ with entries $\frac{1}{\mu_{ij}}$. Rescaling of the linear program analogous for obtaining $(P^{\prime},D^{\prime})$ then yields the following linear programs.
$(\mbox{{Asg}}^{\prime})$
$\text{max}\qquad\sum_{i,j=1}^{n}x_{ij}$
$\text{s.t. }\qquad\sum_{j=1}^{n}\frac{1}{\mu_{ij}}x_{ij}$ $\leq 1\ \forall\,i\in[n]$
$\qquad\sum_{i=1}^{n}\frac{1}{\mu_{ij}}x_{ij}$ $\leq 1\ \forall\,j\in[n]$
$x_{ij}$ $\geq 0\ \forall\,i,j\in[n]$
$(\overline{\mbox{{Asg}}}^{\prime})$
$\text{min}\qquad\sum_{i=1}^{n}u_{i}+\sum_{j=1}^{n}v_{j}$
$\text{s.t. }\qquad\frac{1}{\mu_{ij}}u_{i}+\frac{1}{\mu_{ij}}v_{j}$ $\geq 1\ \forall\,i,j\in[n]$
$u_{i}$ $\geq 0\ \forall\,i\in[n]$
$v_{j}$ $\geq 0\ \forall\,j\in[n]$
The constraint matrix of $(\mbox{{Asg}}^{\prime})$ can now be used to define a zero-sum game, the hide-and-seek game by von Neumann. The game matrix has dimension $2n\times n^{2}$, with only $2n^{2}$ of its $2n^{3}$ entries being nonzero. The nonzero entries correspond exactly to the surplus matrix of the assignment problem.
Recall that we can view the hide-and-seek game as the row and column players picking a vertex and edge of the assignment graph, respectively. The column player tries to hide in the edge and the row player tries to search edges by inspecting end points. Our correspondence explains why the payoffs of the hide-and-seek game are the inverse of the matching surplus of the respective cell. This is due to the rescaling of the linear program to rewrite it in the right form for games. Moreover, we can also revisit von Neumann’s correspondence of maximum weight matchings and strategies of the column player. His main theorem says that the maximin strategies of the column player (von Neumann calls them optimal strategies) are obtained by permutation matrices corresponding to maximum weight matchings. More precisely, if we view a permutation as a bijective function $\sigma\colon[n]\to[n]$, then the maximin strategies are convex combinations of strategies with support determined by permutations $\sigma$ that maximize $\sum_{i=1}^{n}\mu_{i\sigma(i)}$. Within the support of the matching specified by a permutation, the probabilities $p$ must be distributed such that $p_{1}\frac{1}{\mu_{1\sigma(1)}}=\dots=p_{n}\frac{1}{\mu_{n\sigma(n)}}$. [^8] This proportion is captured by von Neumann in the analysis of a so-called $1$-dimensional game that corresponds to the analysis of game play given a fixed permutation matrix (or, equivalently, perfect matching). Hence, it is again the rescaling to obtain $(P^{\prime},D^{\prime})$ that tells how feasible solutions of the primal will be skewed to obtain maximin strategies of the game. Once we consider linear programs of the form $(P^{\prime},D^{\prime})$, the correspondence of optimal primal-dual pairs and maximin strategies requires no more skewing but simply multiplication with a scalar. Put differently, the scaled matrix leads to a uniform scale for the payoffs.
5.2 Production Planning Game
We now provide a more general, natural interpretation of the reduced game for any linear program with positive $b$ and $c$. Such linear programs can be viewed as planning optimal production. A central authority decides how much of each of $n$ goods to produce, where $x_{j}$ is the amount of good $j$ produced. The profit from producing one unit of good $j$ is $c_{j}$. The production of each good consumes certain amounts of $m$ resources representing, e.g., raw materials and electricity, or, more broadly, labor hours and machine capacity. Producing a unit of the $j$th good requires $a_{ij}$ units of the $i$th resource. Finally, we have total availability of $b_{i}$ units for the $i$th resource. The total usage of resources is not allowed to exceed their availability, which is exactly captured by the constraints $Ax\leq b$. Furthermore, it is reasonable to assume positive profits from producing a good as well as a positive availability of each resource. Hence, the production planning scenario is captured by the linear program $(P)$ for positive $b$ and $c$. In many applications, we also have nonnegative entries in $A$. However, negative entries in $A$ allow for modeling further scenarios, such as producing a resource, e.g., electricity, as a byproduct in a production process.
The reduced game now has the following interpretation: the column player mimics the production planner while the row player mimics a resource provider. A pure strategy for the column player is to pick a good while the row player picks a resource. While the production planner wants to gain high production profits, the resource player wants to sell them a scarce resource. Selecting the $j$th good and $i$th resource leads to a payment by the column player to the row player of $\frac{a_{ij}}{b_{i}c_{j}}$. This quantity is the resource share needed to produce a unit of the $j$th good, $\frac{a_{ij}}{b_{i}}$, divided by the profit $c_{j}$ for producing one unit. Together, this is the required resource share per unit profit. If the selected good leads to a small profit or needs a high amount of the resource, i.e., the resource is scarce for this good, then the payment is high. Conversely, needing a small share of the resource or making a lot of profit leads to a small payment, or even a negative payment if the resource is produced as a byproduct indicated by a negative $a_{ij}$. Hence, it makes intuitive sense that the payoff depends on the reciprocals of both the resource availability and the price of the good.
5.3 Playing Linear Programs
We want to broaden the insights of the discussed applications to give an even more general interpretation of our correspondence.
For comparison, we first recall the original construction by Dantzig (1951) as defined in Section 3 given by the game matrix in Equation 1. The skew-symmetric matrix of this game leads to symmetric roles for both players. They both specify strategies of the form $(p,q,t)$, where $q$ corresponds to a solution of the primal linear program $(P)$ and $p$ corresponds to a solution of the dual linear program $(D)$.
By contrast, consider now our construction for positive $b$ and $c$ as detailed in Theorem 4.3. There, the column and row players have different roles corresponding to the primal and dual linear programs, respectively. Moreover, each fixed strategy gives a guarantee to a player, i.e., a minimum reward for the row player against adversarial play of the column player or a maximum payment by the column player against an adversarial row player. These guarantees are obtained by solving $(\mbox{{Row}})$ and $(\mbox{{Col}})$ when fixing strategies $p$ and $q$, respectively. When players play optimal strategies, they get their optimal guarantee, which is exactly the maximin or minimax value. Importantly, such guarantees are obtained for any strategy, not only when players play optimally and they carry information about the underlying LP.
First, if the column player can play a (not necessarily optimal) strategy $q$ that guarantees her negative utility, i.e., for this fixed $q$, the minimum $v$ of $(\mbox{{Col}})$ is negative, then this $q$ gives rise to an unboundedness certificate as in Part 1-(b) of the proof of Theorem 4.3. Interestingly, the unboundedness can not only be certified by an optimal strategy but by any strategy of the column player yielding nonpositive payoff.
Conversely, any strategy that guarantees a positive payoff leads to a feasible solution to the corresponding linear program. Specifically, assume that $\bar{q}$ and $\bar{p}$ are strategies for the column and row players such that $(\mbox{{Col}})$ and $(\mbox{{Row}})$ have positive optimal values $\bar{v}$ and $\bar{u}$ when fixing $\bar{q}$ and $\bar{p}$, respectively. Then, as in Part 1-(b) of the proof of Theorem 4.3, scaling these strategies as $\frac{1}{\bar{v}}\bar{q}$ and $\frac{1}{\bar{u}}\bar{p}$ leads to feasible solutions of $(P^{\prime})$ and $(D^{\prime})$, respectively. Their objective values with respect to $(P^{\prime})$ and $(D^{\prime})$ are $\frac{1}{\bar{v}}$ and $\frac{1}{\bar{u}}$. Once again, this works even for nonoptimal strategies of the reduced games as long as they guarantee a positive payment. However, importantly, such a strategy exists for the row player only if $(D^{\prime})$ admits a feasible solution, and, therefore, $(P^{\prime})$ is not unbounded.
All these considerations give a broader interpretation of the game play as playing a linear program: In the first case, the column player plays an infeasibility certificate. In the second case, the row player can play a feasible strategy, implying that any strategy of the column player has to lead to a positive value of $(\mbox{{Col}})$. Hence, in this case, both players play a feasible solution. This solution might, however, lead to a suboptimal objective value of the LP. In the special case where both play maximin strategies, their play corresponds to playing optimal solutions of the LP, which we can extract alongside the optimal value of the LP.
Notably, as in the case of the assignment problem, the direct correspondence, where solutions only differ by scaling, is obtained between the linear program pair $(P^{\prime},D^{\prime})$ and the zero-sum game based on the constraint matrix of this game. In other words, linear programs of the form $(P^{\prime},D^{\prime})$ are linear programs of games. Von Neumann’s result can therefore be interpreted as the realization that the assignment problem can be written in game form. And more generally, Proposition 4.1 shows that, whenever $b>\mathbf{0}$ and $c>\mathbf{0}$, any linear program pair of the form $(P,D)$ can be written in game form, as well.
6 Relationship with Dantzig’s Reduction
Recall our discussion in Section 1 that Dantzig’s original reduction does not work for the full generality of linear programs. In this section, we explore whether the restrictions that we impose on linear programs suffice to make his reduction work. First, we show that under the assumption $b>\mathbf{0}$ and $c>\mathbf{0}$, his reduction works. Let $(p^{*},q^{*},t^{*})$ be the minimax strategy of the row player for Dantzig’s game as presented in Equation 1. Then, it holds that
$(i)$ $Aq^{*}-bt^{*}\leq\mathbf{0},$
$(ii)$ $-A^{\intercal}p^{*}+ct^{*}\leq\mathbf{0},$
$(iii)$ $b^{\intercal}p^{*}-c^{\intercal}q^{*}\leq\mathbf{0},$
$(iv)$ $\mathbf{1}^{\intercal}p^{*}+\mathbf{1}^{\intercal}q^{*}+t^{*}=1,$
$(v)$ $p^{*}\geq\mathbf{0},\;q^{*}\geq\mathbf{0},\;t^{*}\geq 0,$
$(vi)$ $t^{*}(b^{\intercal}p^{*}-c^{\intercal}q^{*})=0.$
Observe that, as the game is symmetric, $(p^{*},q^{*},t^{*})$ is also a maximin strategy for the column player, and the value of the game is 0.
Proposition 6.1.
The following statements are true:
  1. 1.
    If $t^{*}>0$, then $\left(\frac{1}{t^{*}}q^{*},\frac{1}{t^{*}}p^{*}\right)$ is a primal-dual optimal pair for $(P,D)$.
  2. 2.
    If $t^{*}=0$, then $q^{*}$ is an unboundedness certificate for $(P)$.
Proof.
We prove the statements one by one.
  1. 1.
    We first consider the case $t^{*}>0$. Conditions ($i$), (iiii), and ($v$) imply that $\frac{1}{t^{*}}q^{*}$ and $\frac{1}{t^{*}}p^{*}$ are feasible for $(P)$ and $(D)$, respectively. In addition, by ($vi$), $\frac{1}{t^{*}}\left(b^{\intercal}p^{*}-c^{\intercal}q^{*}\right)=0$, which completes the necessary and sufficient conditions for the optimality of $\frac{1}{t^{*}}q^{*}$ and $\frac{1}{t^{*}}p^{*}$.
  2. 2.
    Now consider the case $t^{*}=0$. Assume for contradiction that $q^{*}=0$. Then, by ($iv$) and ($v$), it holds that $0\neq p^{*}\geq 0$. Hence, since $b>\mathbf{0}$, it follows that $b^{\intercal}p^{*}>0$, contradicting (iiiiii). Thus, $q^{*}\neq 0$ which implies (as $c>\mathbf{0}$) that $c^{\intercal}q^{*}>0$. Moreover, by ($i$) and ($v$), it holds that $Aq^{*}\leq 0$ and $q^{*}\geq 0$. Together, $q^{*}$ is an unboundedness certificate for (P)(P).∎
Notably, the second part of Proposition 6.1 has the precondition $t^{*}=0$ whereas the theorem by Dantzig (1951) had the precondition $b^{\intercal}p^{*}-c^{\intercal}q^{*}<0$. Recall that the “hole” in Dantzig’s proof was due to the fact that it could happen that both $t^{*}=0$ and $b^{\intercal}p^{*}-c^{\intercal}q^{*}=0$. This raises the question whether Proposition 6.1 works because this case never occurs or whether it works because this case can be dealt with successfully (unlike in Dantzig’s reduction). Remarkably, the latter is the case, as our next example shows. Regardless of conditions on $p^{*}$ and $q^{*}$, we obtain an unboundedness certificate whenever $t^{*}=0$.
Example 6.2.
We provide an example illustrating Proposition 6.1 for the case when $t^{*}=0$ and $b^{\intercal}p^{*}-c^{\intercal}q^{*}=0$. Consider the LP given by
$A=\begin{pmatrix}-1&\frac{1}{2}\\ 1&-\frac{1}{2}\end{pmatrix}\text{, }b=\begin{pmatrix}1\\ 1\end{pmatrix}\text{, and }c=\begin{pmatrix}1\\ 1\end{pmatrix}\text{.}$
The matrix of Dantzig’s reduced zero-sum game is given by the matrix
$\begin{pmatrix}0&0&-1&\frac{1}{2}&-1\\ 0&0&1&-\frac{1}{2}&-1\\ 1&-1&0&0&1\\ -\frac{1}{2}&\frac{1}{2}&0&0&1\\ 1&1&-1&-1&0\\ \end{pmatrix}\text{.}$
[…]
In contrast to the case $b>\mathbf{0}$ and $c>\mathbf{0}$ covered by Proposition 6.1, it is possible for Dantzig’s reduction to fail whenever $A\geq\mathbf{0}$.
Example 6.3.
Consider any constraint matrix $A\geq\mathbf{0}$ and consider the linear program pair given by $A$, $b=\mathbf{0}$, and $c=\mathbf{0}$. Dantzig’s reduced game as in Equation 1 for this case is
$\begin{pmatrix}\mathbf{0}&A&\mathbf{0}\\ -A^{\intercal}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}^{\intercal}&\mathbf{0}^{\intercal}&0\\ \end{pmatrix}\text{.}$
[…]
Arguably, the case of $b=c=\mathbf{0}$ in the last example seems somewhat trivial. However, we want to stress that Dantzig’s correspondence can only fail for somewhat trivial cases, namely cases covered by our preprocessing in Algorithm 1. Indeed, instead of using our reduction, we can also use Dantzig’s reduction in Algorithm 1 for tackling LPs with $A\geq\mathbf{0}$. This approach is sound due to Proposition 6.1.
7 Computational Considerations
Up to this point, we have omitted the computational properties of reductions and the number spaces to which they are applicable. In this section, we address these aspects.
7.1 Reasonable Reductions
Typically, in complexity theory, meaningful reductions are required to be polynomial, that is, the effort of the reduction, as well as the effort to recover the solution, is required to be polynomial in the size of the problem. We shall assume for now that the data of the linear programs is rational and discuss implications of our reduction for real input data subsequently. To make the reduction meaningful in the framework of complexity theory and given that any LP problem with rational data is known to be solvable in time polynomial in the size of the problem, we follow Adler (2013) and require reductions to be strongly polynomial, that is, they have to use a polynomial number of arithmetic operations and, if applied to rational input data, polynomial space (Tardos, 1986; Grötschel et al., 1993). In our case, the input data of a linear program depends on its dimensions $m$ and $n$ and the encoding of numbers, hence polynomial means with respect to $m$, $n$, and the logarithm of numbers such as the entries of the constraint matrix. Note that strongly polynomial reductions still permit the occurrence of exponentially large numbers with respect to $m$ and $n$, and hence all the previously existing reductions by Adler (2013), Brooks and Reny (2023), and von Stengel (2023) are strongly polynomial as well. In the context of linear programming, strongly polynomial reductions make a lot of sense since it is a long-standing open problem to decide whether linear programs can be solved in strongly polynomial time (see, e.g., Tardos, 1986; Grötschel et al., 1993, Section 6.6).
Notably, our as well as previous reductions from linear programs to zero-sum games are even log-space reductions, i.e., they only use logarithmic space except for outputting the polynomially large reduced instance (Arora and Barak, 2009). Note that even a log-space reduction can produce exponentially large numbers with respect to input parameters, such as the large numbers in the reductions by Adler (2013), Brooks and Reny (2023), and von Stengel (2023). Their reductions run in log-space because these numbers do not have to be computed with arithmetic operations but can directly be inferred from the input parameters and outputted without computation. [^9] For example, recall the number $m(\sqrt{m}\theta)^{m}+1$ from Footnote 4. In binary representation, this number only uses a polynomial number of bits with respect to the input size. Instead of being computed it can be inferred directly from the input data and simply outputted as a matrix entry of the reduced game. An interesting feature of previous reductions being log-space is that it implies that computing maximin strategies in zero-sum games is P-complete because solving linear programs is P-complete (Dobkin et al., 1979). Apparently, this observation was missed in previous work.
7.2 Reductions for Real Numbers
In the previous subsection, we assumed rational data for the discussion of reasonable reductions. However, in the beginning of Section 2, the problems are presented in terms of real numbers. This presentation is common whenever the mathematics of linear programming problems are discussed. The reason for this is that, as recognized from the early days of linear programming and game theory, all major mathematical theorems (e.g., the LP-Duality Theorem or the Minimax Theorem) are valid for real numbers. [^10] In fact, they are valid for any ordered field (see, e.g., Gale, 1960)). In addition, most related algorithms, such as Fourier-Motzkin elimination or the simplex algorithm, prior to the polynomial-time algorithm for solving LPs by Khachiyan (1979) were developed and presented in terms of real numbers.
Indeed, both the reductions by Dantzig (1951) and von Neumann (1953) only use elementary computations and are applicable for real numbers. However, the corrected extensions of Dantzig’s reduction of Adler (2013), Brooks and Reny (2023), and von Stengel (2023) are only applicable for rational LPs because the constructions require a big number depending on the data. It is currently not known whether this number can be calculated in strongly polynomial time for real-valued LPs. [^11] By contrast, the Turing-reduction by Adler (2013) is applicable for real input but it requires solving up to $n$ LPs.
Not being applicable to real numbers feels like a missing feature of these reductions. LPs and zero-sum games in real numbers are generally relevant for complexity analysis of computational models with real numbers (see, e.g., Blum et al., 1989). In particular, the problem of finding a strongly polynomial algorithm for real-valued LPs is listed as one of the most prominent mathematical challenges for the 21st century (Smale, 1998, Problem 9).
Interestingly, our reduction shares the applicability for real numbers with the original reduction of von Neumann. It thus provides a somewhat more complete equivalence between LPs and zero-sum games, albeit only for positive $b$ and $c$ or nonnegative $A$. Note that only reductions from linear programs to zero-sum games that are applicable to real numbers can be used to translate strongly polynomial algorithms (for real input) for finding maximin strategies for zero-sum games to strongly polynomial algorithms to solve linear programs. In other words, using our technique, any strongly polynomial algorithm for computing maximin strategies for zero-sum games would immediately give rise to a strongly polynomial algorithm for solving linear programs with positive $b$ and $c$ or nonnegative $A$.
8 Discussion
Clearly, the question of generalizing von Neumann’s result (von Neumann, 1953) to an arbitrary primal-dual pair of LPs is an important one. Our work explores an extension based on rescaling, the crucial idea that helps define the payoffs in the hide-and-seek game. Even though this seems only possible if $b>\mathbf{0}$ and $c>\mathbf{0}$, we managed to reduce the case $A\geq\mathbf{0}$ to this case as well. Going further may well require a new approach.
Our paper is first and foremost an attempt at expounding a purely technical result; yet, its historical overtones are hard to miss. In that vein, we would like to end with some observations which may interest students and observers of the history of computing and computability.
A lot has been said about von Neumann’s extraordinary intellectual prowess and his foundational influence on the creation of the digital computer, in particular, his 1945 report on what is known today as the von Neumann architecture (see, e.g., Halmos, 1973; Bhattacharya, 2021). Our study of his masterful work (von Neumann, 1953) has led us to believe that he had an intuitive understanding of the notion of efficient computability, i.e., polynomial-time solvability.
With a view to solving the assignment problem efficiently, von Neumann (1953) gave the following procedure: Reduce the assignment problem to a zero-sum game and approximately solve the latter using methods from Brown and von Neumann (1951), involving solving a sequence of differential equations.
Over the centuries, other mathematicians had also found algorithms that were later seen to be polynomial-time algorithms, e.g., Euclid’s algorithm for computing the greatest common divisor of two numbers. However, von Neumann’s work is qualitatively different. He explicitly reasoned about the time complexity of the method he had proposed, as is evident from reading the first page of the paper (von Neumann, 1953). The page starts by observing that an optimal solution to the assignment problem can be found via an exhaustive search, i.e., by considering all $n!$ permutation matrices, and ends by stating that his procedure gives a “significant approximate solution” in “a moderate power of nn” steps. The page ends with the following striking sentence, which is clearly alluding to the qualitative difference between polynomial and exponential time.
It turns out that this number is a moderate power of $n$, i.e., considerably smaller than the “obvious” estimate $n!$ mentioned earlier.
Additionally:
  1. 1.
    Von Neumann (1953) reduced the assignment problem to the problem of finding a Nash equilibrium in zero-sum games.
  2. 2.
    Today we know that both these problems are in P: the first due to the work of Kuhn (1955) and the second because the problem is equivalent to linear programming, which is in P due to Khachiyan (1979); moreover, the reduction is polynomial time. Recall that the class P was defined more than a decade after von Neumann’s work (Edmonds, 1965).
Hence, von Neumann gave a strongly polynomial-time reduction between two problems in P — over a decade before P was even defined!
Acknowledgements
Most of this research was done while Martin Bullinger was affiliated with the University of Oxford. Martin Bullinger was supported by the AI Programme of The Alan Turing Institute. Vijay V. Vazirani was supported by NSF grant CCF-2230414.

### v3

@@
 game theory (von Neumann and Morgenstern, 1947) but also had far-reaching consequences in other areas.
-In particular, it is equivalent to the Linear Programming (LP) Duality Theorem and it led to Yao’s Lemma (Yao, 1977), a key tool for proving a lower bound on the worst-case performance of a randomized algorithm.
+In particular, it is equivalent to the Linear Programming Duality Theorem and it led to Yao’s Lemma (Yao, 1977), a key tool for proving a lower bound on the worst-case performance of a randomized algorithm.
 The equivalence mentioned above is particularly significant since it connects cornerstone problems of the two fields of game theory and optimization.
 Our paper provides a fresh insight into this equivalence.
-The two directions of this equivalence are qualitatively different: It is straightforward to reduce a zero-sum game to a primal-dual pair of LPs whose optimal solutions yield maximin strategies of the game (Dantzig, 1951).
+The two directions of this equivalence are qualitatively different: It is straightforward to reduce a zero-sum game to a primal-dual pair of linear programs (LPs) whose optimal solutions yield maximin strategies of the game (Dantzig, 1951).
 However, the other direction is much harder, i.e., the task of solving a linear program — or deciding that the primal or the dual program is infeasible — by finding maximin strategies of a corresponding zero-sum game.
@@
 Even though this result stood for over half a century, it is in fact incomplete as observed by Adler (2013): the constructed zero-sum game can have solutions of a particular form which cannot be interpreted as optimal primal-dual pairs or certificates of infeasibility of the primal or the dual LP.
-Later — by adding an additional row to the zero-sum game — this reduction was extended to work in full generality, first by Adler (2013) and later, adding a certificate of infeasibility, by von Stengel (2023) and Brooks and Reny (2023); these are all strongly polynomial (in fact log-space), many-one reductions, see Section 3 for details.
-However, these extensions are valid only for LPs with rational numbers, hence giving the sense that the reduction is still incomplete, since the easy direction is valid for zero-sum games with real numbers (see Section 7.2 for a detailed discussion).
+Later — by adding an additional row to the zero-sum game — this reduction was extended to work in full generality, first by Adler (2013) and later, adding a certificate of infeasibility, by von Stengel (2023) and Brooks and Reny (2023).
 Next, we describe the partial resolution by von Neumann (1953).
-Let $(G,\mu)$ be an instance of the assignment problem on a complete bipartite graph $G$ with $n$ vertices on each side and strictly positive weights $\mu_{ij}$ on all edges $(i,j)$.
+Let $(G,\phi)$ be an instance of the assignment problem on a complete bipartite graph $G$ with $n$ vertices on each side and a strictly positive weight $\phi_{ij}$ on edge $(i,j)$.
 The transformed game matrix $M$ is $2n\times n^{2}$, where the $2n$ rows correspond to the $2n$ vertices of $G$ and the $n^{2}$ columns correspond to the $n^{2}$ edges of $G$.
-Let $l$ represent edge $(i,j)$ of $G$.
-Then the $(k,l)^{th}$ entry of $M$ is defined as follows:
-[…]
+Let $\ell$ represent edge $(i,j)$ of $G$.
+Then the $(k,\ell)^{th}$ entry of $M$ is defined as follows:
+$M(k,\ell)\ \ =\ \begin{cases}1/\phi_{ij}&\text{if }k=i\ \text{ or }k=j,\\ 0&\text{otherwise}.\end{cases}$
 Observe that each column has exactly two nonzero entries.
-The assignment problem asks for a maximum weight matching in bipartite graph $G$ under edge weights $\mu_{ij}$.
+The assignment problem asks for a maximum weight matching in bipartite graph $G$ under edge weights $\phi_{ij}$.
 In the hide-and-seek game, the row player picks one vertex and the column player picks one edge, say $e=(i,j)$, of $G$.
-If the vertex is one of the endpoints of the edge, then the column player pays $\frac{1}{\mu_{ij}}$ to the row player, and otherwise, no payment is made.
+If the vertex is one of the endpoints of the edge, then the column player pays $\frac{1}{\phi_{ij}}$ to the row player, and otherwise, no payment is made.
 The row player attempts to maximize his payoff and the column player minimizes the amount she has to pay.
@@
 The most remarkable step in the construction given above is von Neumann’s choice of the payoff matrix.
-Indeed, up to scaling, this matrix is exactly the constraint matrix of the LP.
+Indeed, up to a change in variables caused by scaling, this matrix is exactly the constraint matrix of the LP.
 He shows that for this payoff matrix, the optimal solutions to the primal and dual LPs yield maximin strategies for the column and row players, respectively.
 But why is the payment the reciprocal of the edge weight?
-Our reduction generalizes the idea of using the scaled constraint matrix as the game matrix, and from this larger perspective, we are able to dispel the mystery behind von Neumann’s reduction, see Section 5.1.
+Our reduction generalizes the idea of using the scaled constraint matrix as the game matrix.
+From this larger perspective, we are able to dispel the mystery behind von Neumann’s reduction, see Section 4.1.
 A drawback of von Neumann’s reduction is that his LPs are highly specialized and have almost a “combinatorial” feel to them: His primal LP is $\max c^{\intercal}x\ \ \text{s.t.
-}Ax\leq\mathbf{1},\ x\geq 0$, with all entries in $c$ being strictly positive and the constraint matrix, $A$, being the incidence matrix of the underlying complete bipartite graph, i.e., all components of this matrix are zeros or ones.
+}Ax\leq\mathbf{1},\ x\geq\mathbf{0}$, with all entries in $c$ being strictly positive and the constraint matrix, $A$, being the incidence matrix of the underlying complete bipartite graph, i.e., all components of this matrix are zeros or ones.
 That still leaves a nagging doubt: can a separation of the roles of the two players be achieved if the LPs are more general, e.g., for a more general constraint vector $b$ or for data with negative entries.
@@
 We leave the open problem of giving a complete generalization.
-We believe this will require a substantially new approach, since we seem to have extended von Neumann’s idea of rescaling as far as was possible.
+We believe this will require a substantially new approach, since we seem to have fully extended von Neumann’s idea of rescaling.
 Our results include:
@@
     Our reduction carves out the exact reason why the hide-and-seek game uses the reciprocals of the assignment values for the payments.
-  4. 4.
-    Our generalization is valid for linear programs in real numbers.
 Finally, since the reduced game always has a well-defined value, achieved by maximin strategies, it offers a new perspective on LP pairs, one of which is unbounded.
-The game-theoretic world allows us to distinguish such LPs; this can be interpreted as a degree of (in)feasibility, see Remark 4.4.
-We end our paper by discussing evidence that von Neumann possessed an implicit understanding of the notion of polynomial time.
+The game-theoretic world allows us to distinguish such LPs; this can be interpreted as a degree of (in)feasibility, see Section 3.1.
 Related Work
 Some previous work has proposed other generalizations of von Neumann’s hide-and-seek game.
-However, in contrast to our paper, none of these generalizations provides a reduction between linear programs and zero-sum games.
+However, in contrast to our paper, none of these provides a reduction between linear programs and zero-sum games.
 First, Marchi and Tarazaga (1981) generalized the hide-and-seek game to higher dimensions.
@@
 We start with some conventions.
-Whenever we write an inequality between vectors, we assume that this inequality holds componentwise.
+Whenever we have an inequality between vectors, we assume that it holds componentwise.
 Moreover, for a vector $v$ and a matrix $M$, we denote by $v^{\intercal}$ and $M^{\intercal}$ its transposed vector and matrix, respectively.
-Finally, we denote by $\mathbf{1}$ and $\mathbf{0}$ the vectors or matrices consisting of $1$- and 0-entries only; their dimension will be clear from the context.
-We now introduce linear programs and zero-sum games, the main objects of our study.
-For text book introductions to these topics, we recommend the books by Schrijver (1998) and Nisan et al.
-(2007).
+Finally, we denote by $\mathbf{1}$ and $\mathbf{0}$ the column vectors or matrices consisting of $1$- and 0-entries only; their dimension will be clear from the context.
+We now introduce linear programs and zero-sum games, the main objects of our study. For text book introductions to these topics, we recommend the books by Schrijver (1998) and Nisan et al. (2007).
 We consider linear programs of the following form.
-Assume that we are given $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$.
+Assume we are given $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$.
 We denote their components by $a_{ij}$, $b_{i}$, and $c_{j}$, respectively.
-These objects specify the linear program $(P)$ together with its dual $(D)$.
+These objects specify the LP $(P)$ together with its dual $(D)$.
 $(P)$ $(D)$
@@
 $\text{s.t. }\qquad Ax$ $\leq b$ $\text{s.t. }\qquad A^{\intercal}y$ $\geq c$
-$x$ $\geq 0$ $y$ $\geq 0$
-We write $(P,D)$ to refer to the pair of linear programs $(P)$ and $(D)$.
-A vector $x\in\mathbb{R}^{n}$ is called (primal) feasible if it satisfies $Ax\leq b$ and $x\geq 0$, i.e., if it satisfies the constraints of $(P)$.
-Similarly, a vector $y\in\mathbb{R}^{m}$ is called (dual) feasible if it satisfies $A^{\intercal}y\geq c$ and $y\geq 0$, i.e., if it satisfies the constraints of $(D)$.
+$x$ $\geq\mathbf{0}$ $y$ $\geq\mathbf{0}$
+We write $(P,D)$ to refer to the pair of LPs $(P)$ and $(D)$.
+A vector $x\in\mathbb{R}^{n}$ is called (primal) feasible if it satisfies $Ax\leq b$ and $x\geq\mathbf{0}$, i.e., if it satisfies the constraints of $(P)$.
+Similarly, a vector $y\in\mathbb{R}^{m}$ is called (dual) feasible if it satisfies $A^{\intercal}y\geq c$ and $y\geq\mathbf{0}$, i.e., if it satisfies the constraints of $(D)$.
 The weak duality theorem states that if $x$ and $y$ are primal feasible and dual feasible, respectively, then $c^{\intercal}x\leq b^{\intercal}y$.
-This means that whenever we find feasible $x$ and $y$ such that $c^{\intercal}x=b^{\intercal}y$, then $x$ is a feasible vector maximizing $c^{\intercal}x$ and $y$ is a feasible vector minimizing $b^{\intercal}y$.
+This implies that whenever we find feasible $x$ and $y$ such that $c^{\intercal}x=b^{\intercal}y$, then $x$ is a feasible vector maximizing $c^{\intercal}x$ and $y$ is a feasible vector minimizing $b^{\intercal}y$.
 In this case, we say that $(x,y)$ an optimal primal-dual pair.
-The strong duality theorem states that whenever $(P)$ and $(D)$ admit feasible points, then they admit an optimal primal-dual pair, and if so, $c^{\intercal}x=b^{\intercal}y$ is called the value of the LP.
-The following characterization of possible outcomes of a pair of primal and dual programs $(P,D)$ is well-known.
+The strong duality theorem states that whenever $(P)$ and $(D)$ admit feasible points, then they admit an optimal primal-dual pair, and if so, $c^{\intercal}x=b^{\intercal}y$; this is called the value of the LP.
+The following characterization of possible outcomes of a pair of primal and dual LPs $(P,D)$ is well-known.
 Proposition 2.1.
@@
     In this case $x^{*}$ and $y^{*}$ are optimal solutions to $(P)$ and $(D)$ respectively.
+  2. 2.
     There exists $z^{*}$ such that
@@
     In this case $(P)$ is infeasible. We call $z^{*}$ an infeasibility certificate for $(P)$.
-  2. 2.
+  3. 3.
     There exists $w^{*}$ such that
@@
 We make a few important remarks.
-In order to certify $(P)$ to be unbounded, we need to show, in addition to the unboundedness certificate, a feasible solution to $(P)$ (i.e., a vector $x$ such that $Ax\leq b$, $x\geq\mathbf{0})$.
-If $z^{*}$ is an infeasibility certificate for $(P)$, then $z^{*}$ is an unboundedness certificate for $(D)$.
-If $w^{*}$ is an unboundedness certificate for $(P)$, then $w^{*}$ is an infeasibility certificate for $(D)$.
-Cases (2) and (3) in Proposition 2.1 establish that the linear programs have no optimal solutions by certifying that either $(P)$ or $(D)$ have no feasible solution.
+$(i)$ In order to certify $(P)$ to be unbounded, we need to show, in addition to the unboundedness certificate, a feasible solution to $(P)$ (i.e., a vector $x$ such that $Ax\leq b$, $x\geq\mathbf{0})$.
+$(ii)$ If $z^{*}$ is an infeasibility certificate for $(P)$, then $z^{*}$ is an unboundedness certificate for $(D)$.
+Similarly, if $w^{*}$ is an unboundedness certificate for $(P)$, then $w^{*}$ is an infeasibility certificate for $(D)$.
+$(iii)$ Cases (2) and (3) in Proposition 2.1 establish that the LPs have no optimal solutions by certifying that either $(P)$ or $(D)$ are infeasible.
 Our second important concept is zero-sum games.
 A zero-sum game is specified by a matrix $M=(m_{ij})\in\mathbb{R}^{m\times n}$.
-There are two players, a row player and a column player; for ease of notation, we will assume that the first is male and the second is female.
+There are two players, a row player and a column player; for ease of description, we will assume that the first is male and the second is female.
 The row player’s action is to select a row $i$ with $1\leq i\leq m$ and the column player’s action is to select a column $j$ with $1\leq j\leq n$.
@@
 Hence, whereas the row player’s goal is to maximize the value of the matrix entry chosen, the column player’s goal is to minimize it.
-The players are allowed to randomize their actions.
+Players are allowed to randomize their actions.
 A mixed strategy is a probability distribution over actions.
@@
 The unique payments determined by these strategies, i.e., $u^{*}=\max_{p}\min_{q}p^{\intercal}Mq$ and $v^{*}=\min_{q}\max_{p}p^{\intercal}Mq$ are called the maximin value and minimax value of the game, respectively.
-The famous Minimax Theorem by von Neumann (1928) states that
+The famous Minimax Theorem of von Neumann (1928) states that
 $u^{*}=\max_{p}\min_{q}p^{\intercal}Mq=\min_{q}\max_{p}p^{\intercal}Mq=v^{*}\text{,}$
 i.e., the maximin value of the game coincides with its minimax value.
-Therefore, this quantity is simply called the value of the game.
+Therefore, this quantity is called the value of the game.
 The goal of solving a zero-sum game is to compute a pair of maximin strategies.
 Note that such strategies are mutual best responses, and therefore constitute a Nash equilibrium.
-It is well known that maximin strategies as well as the value of the game can be computed by the following linear programs; this is the easy direction of the correspondence between linear programs and zero-sum games (see, e.g., von Stengel, 2023, Equations (7) and (8)):
+It is well known that maximin strategies as well as the value of the game can be computed by the following LPs; this is the easy direction of the correspondence between LPs and zero-sum games (see, e.g., von Stengel, 2023, Equations (7) and (8)):
 $($ $\mbox{{Row}})$ $~($ $\mbox{{Col}})$
 max $u$ min $v$
-s.t. $M^{\intercal}p\geq\mathbf{1}u$ s.t. $Mq\leq\mathbf{1}v$
+s.t. $M^{\intercal}p\geq u\mathbf{1}$ s.t. $Mq\leq v\mathbf{1}$
 $\mathbf{1}^{\intercal}p=1,\;p\geq\mathbf{0}$ $\mathbf{1}^{\intercal}q=1,\;q\geq\mathbf{0}$
-Note that $(\mbox{{Row}},\mbox{{Col}})$ are a primal-dual pair.
+Note that $(\mbox{{Row}},\mbox{{Col}})$ are a primal-dual LP pair.
 The optimal value of the linear program $(\mbox{{Row}})$ is the maximin value corresponding to the row player.
@@
 Similarly, $(\mbox{{Col}})$ determines the minimax value and a minimax strategy of the column player.
-3 Reducing Linear Programs to Zero-Sum Games
-In this section, we lay the groundwork for performing the difficult direction of the correspondence, i.e., reductions from LPs to zero-sum games.
-We start by reviewing the ideas behind the original reduction by Dantzig (1951).
-In his seminal paper already discussed in the introduction, Dantzig (1951) introduced a reduction from LPs to zero-sum games by using a skew-symmetric game matrix.
-More precisely, given the primal LP $(P)$, he considers the zero-sum game with payoff matrix [^3] Dantzig (1951) attributes the idea for this matrix to Tucker.
-$\begin{pmatrix}\vskip 6.0pt plus 2.0pt minus 2.0pt\mathbf{0}&A&-b\\ -A^{\intercal}&\mathbf{0}&c\\ b^{\intercal}&-c^{\intercal}&0\\ \end{pmatrix}\text{.}$ (1)
-As this matrix is skew-symmetric (so it is a symmetric game), the value of the game is 0.
-Denoting a maximin strategy as $(p^{*},q^{*},t^{*})$, the key result of the paper by Dantzig (1951) states that
-  1. 1.
+3 Extending Von Neumann’s Reduction
+In this section, we provide our extensions of von Neumann’s reduction.
+First, we consider the case where $b>\mathbf{0}$ and $c>\mathbf{0}$, i.e., both vectors are strictly positive.
+Subsequently, we will show how this enables us to tackle the case that the constraint matrix is nonnegative, i.e., $A\geq\mathbf{0}$.
+3.1 Positive Objective and Constraint Vectors
+The key technical step of our reduction is a rescaling of the LPs $(P)$ and $(D)$; this is where we require $b>\mathbf{0}$ and $c>\mathbf{0}$.
+Consider the diagonal matrices $B\in\mathbb{R}^{m\times m}$ and $C\in\mathbb{R}^{n\times n}$ defined by
 […]
-  2. 2.
-    If $b^{\intercal}p^{*}-c^{\intercal}q^{*}<0$, then either $z^{*}=p^{*}$ satisfies (2) or $w^{*}=q^{*}$ satisfies (3) in Proposition 2.1.
-Note that since $(p^{*},q^{*},t^{*})$ is a maximin strategy for a game of value 0, it holds that $t^{*}(b^{\intercal}p^{*}-c^{\intercal}q^{*})=0$.
-The goal of solving $(P)$ is achieved, except for the case $t^{*}=b^{\intercal}p^{*}-c^{\intercal}q^{*}=0$.
-Notably, this “hole” in the reduction was clearly recognized by Dantzig (1951).
-Incredibly, Dantzig’s reduction was nonetheless accepted and ascribed as complete for fifty years.
-Only then, a series of papers have worked towards a complete reduction.
-First, Adler (2013) modified Dantzig’s reduction and proposed a many-one reduction for the case that $A$, $b$, and $c$ are rational numbers.
-In the sequel, both Brooks and Reny (2023) and von Stengel (2023) noted that the output of Adler’s reduction was either a pair $(x^{*},y^{*})$ satisfying (1) in Proposition 2.1 or an assurance that there exists either $z^{*}$ or $w^{*}$ (possibly both) satisfying (2) or (3) in Proposition 2.1.
-Both papers introduce modifications adding a specific certificate of infeasibility that allows for solving $(P)$ as stated in Proposition 2.1.
-The key idea of these papers is to modify the reduced game by Dantzig: Adler (2013) and von Stengel (2023) add an additional row while Brooks and Reny (2023) transform the game with elementary column operations.
-These reductions have in common the fact that the constraint matrix or vectors in the reduced games need large numbers, exponential in $m$ or nn.
-[^4] As an example, in the reduction by Adler (2013), the additional entry of the constraint vector is estimated as a number that is at least $m(\sqrt{m}\theta)^{m}+1$ where $\theta$ is the absolute value of the largest number contained in $A$ and $b$.
-A second way to solve linear programs explored by Adler (2013) is to use the original reduction by Dantzig and sequentially eliminate variables until the solution is in the right form.
-While this avoids large numbers, and is applicable to real numbers, it requires solving more than one zero-sum game.
-In other words, it is a Turing reduction rather than a many-one reduction.
-[^5] We refer to Section 7 for an in-depth discussion of the computational properties of our reduction.
-4 Extending Von Neumann’s Reduction
-In this section, we provide our extension of von Neumann’s reduction.
-First, we consider the case that the vectors $b>\mathbf{0}$ and $c>\mathbf{0}$, i.e., both are strictly positive.
-Next we will show how this enables us to tackle the case that the constraint matrix is nonnegative, i.e., $A\geq\mathbf{0}$.
-4.1 Positive Objective and Constraint Vectors
-The key technical step of our reduction is a rescaling of the linear programs $(P)$ and $(D)$; this is where we require $b>\mathbf{0}$ and $c>\mathbf{0}$, i.e., both vectors need to be strictly positive.
-We consider the diagonal matrices $B\in\mathbb{R}^{m\times m}$ and $C\in\mathbb{R}^{n\times n}$ defined by
-[…]
-We define the matrix $M=BAC\in\mathbb{R}^{m\times n}$ (its name already suggests that this will be the payoff matrix of our reduced zero-sum game) and a related pair of linear programs as follows.
+We define the matrix $M=BAC\in\mathbb{R}^{m\times n}$ (its name already suggests that this will be the payoff matrix of our reduced zero-sum game) and a related pair of LPs as follows.
 $(P^{\prime})$ $(D^{\prime})$
@@
 $\xi$ $\geq\mathbf{0}$ $\eta$ $\geq\mathbf{0}$
-The relationship with the original linear programs is captured as follows.
-Proposition 4.1.
-Assume that $b>\mathbf{0}$ and $c>\mathbf{0}$. The following statements are true:
-  1. 1.
-    1. (a)
-      $\xi$ is feasible for $(P^{\prime})$ if and only if $C\xi$ is feasible for $(P)$,
-    2. (b)
-      $\eta$ is feasible for $(D^{\prime})$ if and only if $B\eta$ is feasible for $(D)$.
-  2. 2.
-    $(\xi^{*},\eta^{*})$ is an optimal primal-dual pair for $(P^{\prime})$ and $(D^{\prime})$ if and only if $(C\xi^{*},B\eta^{*})$ is an optimal primal-dual pair for $(P)$ and $(D)$.
-  3. 3.
+It is easy to see that this change of variables by scaling merely leads to a rescaling in the solutions of the original pair of LPs:
+  1. (i)
+    $\xi$ (resp., $\eta$) is feasible for $(P^{\prime})$ (resp., $(D^{\prime})$) if and only if $C\xi$ (resp., $B\eta$) is feasible for $(P)$ (resp., $(D)$)
+  2. (ii)
+    $(\xi^{*},\eta^{*})$ is an optimal primal-dual pair for $(P^{\prime},D^{\prime})$ if and only if $(C\xi^{*},B\eta^{*})$ is an optimal primal-dual pair for $(P,D)$.
+  3. (iii)
     $\bar{\xi}$ is an unboundedness certificate for $(P^{\prime})$ if and only if $C\bar{\xi}$ is an unboundedness certificate for $(P)$.
-Proof.
-We prove the statements one by one.
-  1. 1.
-    Clearly, $\xi\geq 0$ if and only if $C\xi\geq 0$. Moreover, $M\xi\leq\mathbf{1}$ if and only if $A(C\xi)=B^{-1}M\xi\leq B^{-1}\mathbf{1}=b$. This implies (a). The proof of (b) is analogous.
-  2. 2.
-    From ($a$) and ($b$) we have that $\xi,\eta$ are feasible for $(P^{\prime}),(D^{\prime})$ if and only if $C\xi,B\eta$ are feasible for $(P),(D)$, respectively.
-In addition, $c^{\intercal}(C\xi)=\mathbf{1}^{\intercal}\xi$ and $b^{\intercal}(B\eta)=\mathbf{1}^{\intercal}\eta$.
-Applying proposition 2.1, completes the proof.
-  3. 3.
-    This follows from (1) and the observation that, because $c>\mathbf{0}$, $\mathbf{1}^{\intercal}\xi>0$ if and only if $\mathbf{1}^{\intercal}C\xi>0$.
-Hence, $\bar{\xi}\in\{\xi\;|\;M\xi\leq\mathbf{0},\;\xi\geq\mathbf{0},\;\mathbf{1}^{\intercal}\xi>0\}$ if and only if $C\bar{\xi}\in\{x\;|\;Ax\leq\mathbf{0},\;x\geq\mathbf{0},\;\mathbf{1}^{\intercal}x>0\}$.∎
-Remark 4.2.
-Proposition 4.1 covers cases (1) and (2) of Proposition 2.1, i.e., there exists an optimal primal-dual pair of solutions for $(P)$ or that $(P)$ is unbounded.
-Case (LABEL:it:inf), when $(P)$ and $(P^{\prime})$ are infeasible, never occurs under the assumption that $b>\mathbf{0}$, since in this case, $x=\mathbf{0}$ is a primal feasible solution.
-Let $(G)$ be the zero-sum game with payoff matrix $M$ where the row-player wants to maximize his payoff and the column player wants to minimize her payoff.
-Recall that the maximin and minimax values of this game correspond to the optimal values of the linear programs introduced in Section 2:
-$($ $\mbox{{Row}})$ $~($ $\mbox{{Col}})$
-max $u$ min $v$
-s.t. $M^{\intercal}p\geq\mathbf{1}u$ s.t. $Mq\leq\mathbf{1}v$
-$\mathbf{1}^{\intercal}p=1,\;p\geq 0$ $\mathbf{1}^{\intercal}q=1,\;q\geq 0$
+We remark that $(P)$ and $(P^{\prime})$ are never infeasible when $b>\mathbf{0}$, since in this case, $x=\mathbf{0}$ is a feasible solution.
+Now, let $(G)$ be the zero-sum game with payoff matrix $M$ where the row-player wants to maximize his payoff and the column player wants to minimize her payoff.
+Recall that the maximin and minimax values of this game correspond to the optimal values of the LPs $(\mbox{{Row}})$ and $(\mbox{{Col}})$ defined in Section 2.
 We have the following theorem.
+Note that while it establishes a correspondence between zero-sum games and LP pairs of the type $(P^{\prime},D^{\prime})$, it applies to LP pairs of the type $(P,D)$ whenever $b>\mathbf{0}$ and $c>\mathbf{0}$, i.e., the rescaling can be performed.
 Theorem 3.1.
-Assume that $b>\mathbf{0}$ and $c>\mathbf{0}$.
   1. 1.
@@
     1. (a)
-      If $v^{*}>0$, then $\left(\frac{1}{v^{*}}q^{*},\frac{1}{v^{*}}p^{*}\right)$ is an optimal primal-dual pair for $(P^{\prime})$ and $(D^{\prime})$.
+      If $v^{*}>0$, then $\left(\frac{1}{v^{*}}q^{*},\frac{1}{v^{*}}p^{*}\right)$ is an optimal primal-dual pair for $(P^{\prime},D^{\prime})$.
     2. (b)
@@
   2. 2.
-    Let $v^{*}$ be the value of the game $(G)$.
+    Let $v^{*}$ be the value of the zero-sum game $(G)$.
     1. (a)
@@
     2. (b)
-      If $w^{*}$ is an unboundedness certificate of $(P^{\prime})$, then $v^{*}\leq 0$.
+      If $(P^{\prime})$ is unbounded, then $v^{*}\leq 0$.
 Proof.
 We prove the statements one by one.
-  1. 1-(a)
+  1. 1-(a).
     The feasibility of $\left(\frac{1}{v^{*}}q^{*},\frac{1}{v^{*}}p^{*}\right)$ for $(P^{\prime})$ and $(D^{\prime})$ is readily verified by the feasibility of $(p^{*},q^{*})$ for $(\mbox{{Col}})$ and $(\mbox{{Row}})$.
 In addition, $\mathbf{1}^{\intercal}p^{*}=\mathbf{1}^{\intercal}q^{*}=1$ implies $\mathbf{1}^{\intercal}\frac{1}{v^{*}}q^{*}=\mathbf{1}^{\intercal}\frac{1}{v^{*}}p^{*}$ which completes the proof.
-  2. 1-(b)
+  2. 1-(b).
     Suppose that $v^{*}\leq 0$.
-Then, $Mq^{*}\leq\mathbf{1}v^{*}\leq\mathbf{0}$, $\mathbf{1}^{\intercal}q^{*}=1>0$, and $q^{*}\geq\mathbf{0}$.
+Then, $Mq^{*}\leq v^{*}\mathbf{1}\leq\mathbf{0}$, $\mathbf{1}^{\intercal}q^{*}=1>0$, and $q^{*}\geq\mathbf{0}$.
 Hence, $q^{*}$ is an unboundedness certificate for $(P^{\prime})$.
-  3. 2-(a)
+  3. 2-(a).
     First, note that the optimal objective value of $(P^{\prime})$ is always strictly larger than 0, and, therefore, $\mathbf{1}^{\intercal}\xi^{*}>0$.
 Define $\hat{v}:=\frac{1}{\mathbf{1}^{\intercal}\xi^{*}}>0$.
-We have that $\mathbf{1}^{\intercal}(\hat{v}\eta^{*})=\mathbf{1}^{\intercal}(\hat{v}\xi^{*})=1$, $\hat{v}\eta^{*}\geq 0$, and $\hat{v}\xi^{*}\geq 0$, and, therefore, $\hat{v}\eta^{*}$ and $\hat{v}\xi^{*}$ are valid strategies.
+We have that $\mathbf{1}^{\intercal}(\hat{v}\eta^{*})=\mathbf{1}^{\intercal}(\hat{v}\xi^{*})=1$, $\hat{v}\eta^{*}\geq\mathbf{0}$, and $\hat{v}\xi^{*}\geq\mathbf{0}$, and, therefore, $\hat{v}\eta^{*}$ and $\hat{v}\xi^{*}$ are valid strategies.
     Next, we show that $\hat{v}\eta^{*}$ and $\hat{v}\xi^{*}$ are mutual best responses. Consider any strategy $q$ of the column player. By dual feasibility, it holds that
@@
     Finally, combining Equation 1 for $q=\hat{v}\xi^{*}$ and Equation 2, we know that $(\hat{v}\eta^{*})^{\intercal}M(\hat{v}\xi^{*})=\hat{v}$. Hence, $v^{*}=\hat{v}$ is the value of the game $(G)$.
-  4. 2-(b)
+  4. 2-(b).
     Assume that $w^{*}$ is an unboundedness certificate for $(P^{\prime})$, i.e., $Mw^{*}\leq\mathbf{0}$, $w^{*}\geq\mathbf{0}$, and $\mathbf{1}^{\intercal}w^{*}>0$.
@@
 We claim that $(q,v)$ is feasible for $(\mbox{{Col}})$.
-    Indeed, since $w^{*}\geq 0$, it holds that $\alpha w^{*}\geq 0$.
+    Indeed, since $w^{*}\geq\mathbf{0}$, it holds that $\alpha w^{*}\geq\mathbf{0}$.
 In addition, $\mathbf{1}^{\intercal}q=\alpha\mathbf{1}^{\intercal}w^{*}=1$.
-Moreover, $Mq=\alpha Mw^{*}\leq\alpha\mathbf{0}=\mathbf{1}v$.
+Moreover, $Mq=\alpha Mw^{*}\leq\alpha\mathbf{0}=v\mathbf{1}$.
     Hence, $(q,v)$ is feasible for $(\mbox{{Col}})$ and therefore $v^{*}\leq 0$. ∎
+We conclude this subsection with remarks about the interpretation of our reduction.
+Linear Programs in Game Form
 An appealing aspect of our reduction is that it uses a scaled version of the constraint matrix, $A$, as the game matrix.
@@
 In addition to being natural, this approach strengthens the connection between the original linear program and the reduced game beyond what was achieved by Dantzig and his successors.
-In particular, the games by Dantzig and Adler are symmetric and have value zero, i.e., the value of the game has no information.
+Once the rescaling is applied, we obtain a correspondence by directly using the constraint matrix as the game matrix of a zero-sum game.
+In other words, LPs of the form $(P^{\prime},D^{\prime})$ can be interpreted as linear programs in game form.
+Hence, von Neumann’s hide-and-seek game results from the realization that the assignment problem can be written in game form.
+And more generally, we observe that any linear program pair of the form $(P,D)$ with $b>\mathbf{0}$ and $c>\mathbf{0}$ can be transformed to game form, as well.
+Unlike in previous reductions, the column and row players in our reduction have different roles corresponding to the primal and dual linear programs, respectively.
+Each fixed strategy (including a nonoptimal one) gives a guarantee to a player, i.e., a minimum reward for the row player against adversarial play of the column player or a maximum payment by the column player against an adversarial row player.
+These guarantees are obtained by solving $(\mbox{{Row}})$ and $(\mbox{{Col}})$ after fixing strategies $p$ and $q$, respectively.
+When players play optimal strategies, they get their optimal guarantee, which is exactly the maximin or minimax value.
+In fact, one can view playing the game of an LP in game form as playing the linear program.
+Any strategy guaranteeing the column player nonpositive payoff (not only the maximin strategy) gives rise to an unboundedness certificate as in Part 1-(b) of the proof of Theorem 3.1.
+Moreover, strategies guaranteeing the row player a positive payoff correspond to dual feasible points.
+Once again, this works even for nonoptimal strategies, but can then lead to feasible points with a suboptimal objective value.
+Hence, players compete by proposing (good) feasible points of their respective LPs, whenever this is possible.
+This can be viewed as playing an LP.
+The reduced zero-sum game yields a reasonable outcome of playing, even for suboptimal or infeasible strategies.
+In the special case, where both players play maximin strategies, we obtain an optimal primal-dual pair.
+Value of Game as Degree of Infeasibility
+The games by Dantzig and Adler are symmetric and have value zero, i.e., the value of the game has no information.
 [^3] While we regard the nonsymmetry of our reduced games as beneficial for interpreting the game, we acknowledge that reductions producing symmetric games have other, computational advantages.
@@
 (1950) imply that algorithms for solving symmetric games can be applied to nonsymmetric games without significant further computational effort.
-The value of the reduced games by von Stengel as well as Brooks and Reny are zero if and only if there exists an optimal solution to the original LP.
-In contrast, the value of our reduced game has valuable information: it allows us to decide whether the LP admits an optimal solution and, if this is the case, it is the reciprocal of its optimal value.
-We believe this is a potent fact which may find a suitable application.
-We conclude this subsection with a remark about how to interpret the value of the reduced game as a degree of (in)feasibility.
-Remark 4.4.
-The primal-dual pair $(\mbox{{Row}},\mbox{{Col}})$ always has primal and dual feasible solutions, whereas $(D^{\prime})$ might be infeasible.
-Recall from Remark 2.2 that an infeasibility certificate for $(D^{\prime})$ is equivalent to an unboundedness certificate for $(P^{\prime})$, i.e., we ask whether there exists $w^{*}$ with $Mw^{*}\leq\mathbf{0}$, $w^{*}\leq 0$, and $\mathbf{1}^{\intercal}w^{*}>0$.
+Moreover, the value of the reduced games by von Stengel as well as Brooks and Reny are zero if and only if there exists an optimal solution to the original LP.
+In contrast, the value of our reduced game has valuable information: first of all, it allows us to decide whether the LP admits an optimal solution and, if so, it is the reciprocal of its optimal value.
+Additionally, the primal-dual pair $(\mbox{{Row}},\mbox{{Col}})$ always has primal and dual feasible solutions, whereas $(D^{\prime})$ might be infeasible.
+Recall from Remark 2.2 that an infeasibility certificate for $(D^{\prime})$ is equivalent to an unboundedness certificate for $(P^{\prime})$, i.e., we ask whether there exists $w^{*}$ with $Mw^{*}\leq\mathbf{0}$, $w^{*}\geq\mathbf{0}$, and $\mathbf{1}^{\intercal}w^{*}>0$.
 By scaling $w^{*}$, the last condition can equivalently be written as $\mathbf{1}^{\intercal}w^{*}=1$.
@@
 3.2 Nonnegative Constraint Matrix
-Finally, we show that the correspondence described in Theorem 4.3 is applicable in a different setting as well.
-We demonstrate how to use the developed theory to deal with linear program pairs $(P,D)$ when $A\geq\mathbf{0}$ but $b$ and $c$ are not constrained to be positive.
-The intuitive idea is to preprocess the linear program to satisfy the conditions that $b$ and $c$ are positive: If $b_{i}<0$, then the nonnegativity of $A$ immediately implies infeasibility.
+Finally, we show that the developed theory in Section 3.1 also allows us to deal with linear program pairs $(P,D)$ when $A\geq\mathbf{0}$ but $b$ and $c$ are not constrained to be positive.
+The intuitive idea is to preprocess the LP to satisfy the conditions that $b$ and $c$ are positive: If $b_{i}<0$, then the nonnegativity of $A$ immediately implies infeasibility.
 Otherwise, we want to delete all columns and rows that correspond to nonpositive entries of $b$ and $c$.
@@
 Similarly, $c_{j}\leq 0$ requires $x_{j}=0$ and we can omit such columns to transition to a smaller LP.
-We thus reach an LP of the original form, for which we can apply the correspondence of Section 4.1.
+We thus reach a reduced LP with $b>\mathbf{0}$ and $c>\mathbf{0}$, for which we can apply the correspondence of Section 3.1.
 We now define sets $I_{0}$, $J_{0}$, and $J_{-}$ of indices of rows and columns to be deleted according to the above discussion.
@@
 $\text{s.t. }\qquad\hat{A}x$ $\leq\hat{b}$ $\text{s.t. }\qquad\hat{A}^{\intercal}y$ $\geq\hat{c}$
-$x$ $\geq 0$ $y$ $\geq 0$
+$x$ $\geq\mathbf{0}$ $y$ $\geq\mathbf{0}$
 […]
 […]
-We collect the key insights for the correctness of Algorithm 1 in the following theorem.
 Input: Linear program pair specified by $A\geq\mathbf{0}$, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$
@@
     $(P)$ is feasible if and only if $b\geq\mathbf{0}$. If there exists $i\in[m]$ such that $b_{i}<0$, then Algorithm 1 returns an infeasibility certificate for $(P)$ in algorithm 1.
+Now suppose $b\geq\mathbf{0}$.
   1. 2.
-    Suppose that $b\geq\mathbf{0}$.
-    1. (a)
     $(P)$ is unbounded if and only if there exists $j\in[n]\setminus J_{-}$ such that $a_{ij}=0$ for all $i\in[m]$.
 In this case, Algorithm 1 returns an unboundedness certificate for $(P)$ in algorithm 1.
+  2. 3.
+    Otherwise $(P,D)$ admits an optimal primal-dual solution. There are two cases for Algorithm 1:
+    1. (a)
+      If $I_{0}=[m]$ or $J_{0}\cup J_{-}=[n]$, then Algorithm 1 returns an optimal primal-dual pair for $(P,D)$ in algorithm 1.
     2. (b)
-      If $I_{0}=[m]$ or $J_{0}\cup J_{-}=[n]$, then Algorithm 1 returns an optimal primal-dual pair for $(P,D)$ in algorithm 1.
-  3. 3.
 […]
@@
 We prove the statements one by one.
-1.
+  1. 1.
     If $b\geq\mathbf{0}$, then $x=\mathbf{0}$ is a feasible solution of $(P)$.
@@
 Hence, $z^{*}$ is an infeasibility certificate for $(P)$.
-2-(a)
+  2. 2.
     Assume that there exists $j\in[n]\setminus J_{-}$ such that $a_{ij}=0$ for all $i\in[m]$.
@@
 Since $w^{*}_{j^{*}}>0$, it follows that $A_{j^{*}}\leq\mathbf{0}$ and therefore $A_{j^{*}}=\mathbf{0}$ by the nonnegativity of $A$.
-2-(b)
+  3. 3-(a).
     Consider the pair $(x^{*},y^{*})$ returned in algorithm 1.
 Clearly $x^{*}=\mathbf{0}$ is primal feasible.
-Moreover, $y^{*}\geq 0$.
+Moreover, $y^{*}\geq\mathbf{0}$.
 Now let $j\in[n]$.
@@
 Together, $(x^{*},y^{*})$ is an optimal primal-dual pair.
-3.
+  4. 3-(b).
     We will first prove the existence of feasible solutions to $\hat{P}$ and $\hat{D}$. Clearly $\hat{x}=\mathbf{0}$ is feasible for (P^)($\hat${P}) because $\hat{b}>\mathbf{0}$.
-Moreover, since $\hat{b}\geq 0$, we can apply part 2-(a) to conclude that (P^)($\hat${P}) is bounded if $\hat{A}$ does not contain 0-columns (which is a stronger condition than only requiring certain nonzero columns).
+    Moreover, since $\hat{b}\geq\mathbf{0}$, we can apply part 2-(a) to conclude that (P^)($\hat${P}) is bounded if $\hat{A}$ does not contain 0-columns (which is a stronger condition than only requiring certain nonzero columns).
 Let $j\in[n]\setminus(J_{0}\cup J_{-})$.
@@
 Algorithm 1 solves linear programs with $A\geq\mathbf{0}$.
-5 Application and Interpretation
-In this section, we want to demonstrate the applicability of our reductions.
-We start by deepening our understanding of von Neumann’s hide-and-seek game in the context of our more general reduction.
-We then discuss playing the reduced game for any LP with positive $b$ and $c$.
-Finally, we provide an interpretation of the reduced game in relationship to its original LP, where we interpret the players’ strategies as “playing” the primal and dual LP.
-5.1 Von Neumann’s Correspondence for the Assignment Problem
-Originally, von Neumann had designed his reduction for the special case where the source LP corresponds to the (fractional) assignment problem.
-The assignment problem consists of $n$ workers and $n$ jobs and it is assumed that assigning worker $i$ to job $j$ creates a positive surplus $\mu_{ij}>0$.
-A fractional assignment of workers to jobs is a nonnegative matrix $(x_{ij})$, where
-  • for all $1\leq i\leq n$, it holds that $\sum_{j=1}^{n}x_{ij}\leq 1$, and
-  • for all $1\leq j\leq n$, it holds that $\sum_{i=1}^{n}x_{ij}\leq 1$.
-The constraints ensure that each worker is assigned no more than the fractions equivalent to one job, and no job is assigned more than one worker.
-The goal is to find a fractional assignment that maximizes the expected surplus.
-[^7] Due to the famous algorithm by Birkhoff (1946), an optimal fractional assignment can be decomposed into perfect matchings of the same value.
-Hence, finding an optimal fractional assignment is sufficient to find an optimal integral assignment.
-In fact, the paper by von Neumann (1953) is more famous as a reference for the Birkhoff-von Neumann Decomposition Theorem of bistochastic matrices rather than for introducing the hide-and-seek game.
-This is equivalent to solving the linear program $(\mbox{{Asg}})$ whose dual is given by $(\overline{\mbox{{Asg}}})$.
-$(\mbox{{Asg}})$
-$\text{max}\qquad\sum_{i,j=1}^{n}\mu_{ij}x_{ij}$
-$\text{s.t. }\qquad\sum_{j=1}^{n}x_{ij}$ $\leq 1\ \forall\,i\in[n]$
-$\qquad\sum_{i=1}^{n}x_{ij}$ $\leq 1\ \forall\,j\in[n]$
-$x_{ij}$ $\geq 0\ \forall\,i,j\in[n]$
-$(\overline{\mbox{{Asg}}})$
-$\text{min}\qquad\sum_{i=1}^{n}u_{i}+\sum_{j=1}^{n}v_{j}$
-$\text{s.t. }\qquad u_{i}+v_{j}$ $\geq\mu_{ij}\ \forall\,i,j\in[n]$
-$u_{i}$ $\geq 0\ \forall\,i\in[n]$
-$v_{j}$ $\geq 0\ \forall\,j\in[n]$
-(Asg) is a linear program of the form $(P)$ where the objective vector is strictly positive (containing the entries $\mu_{ij}$), the constraint matrix contains only zeros and ones, and the right hand side is the all-ones vector.
-Its dual contains variables $u_{i}$ and $v_{j}$ corresponding to workers and jobs of the assignment problem.
-Now, matrix $B$ for $(\mbox{{Asg}})$ is the identity matrix, and matrix $C$ is the diagonal matrix of dimension $n^{2}\times n^{2}$ with entries $\frac{1}{\mu_{ij}}$.
-Rescaling of the linear program analogous for obtaining $(P^{\prime},D^{\prime})$ then yields the following linear programs.
-$(\mbox{{Asg}}^{\prime})$
-$\text{max}\qquad\sum_{i,j=1}^{n}x_{ij}$
-$\text{s.t. }\qquad\sum_{j=1}^{n}\frac{1}{\mu_{ij}}x_{ij}$ $\leq 1\ \forall\,i\in[n]$
-$\qquad\sum_{i=1}^{n}\frac{1}{\mu_{ij}}x_{ij}$ $\leq 1\ \forall\,j\in[n]$
-$x_{ij}$ $\geq 0\ \forall\,i,j\in[n]$
-$(\overline{\mbox{{Asg}}}^{\prime})$
-$\text{min}\qquad\sum_{i=1}^{n}u_{i}+\sum_{j=1}^{n}v_{j}$
-$\text{s.t. }\qquad\frac{1}{\mu_{ij}}u_{i}+\frac{1}{\mu_{ij}}v_{j}$ $\geq 1\ \forall\,i,j\in[n]$
-$u_{i}$ $\geq 0\ \forall\,i\in[n]$
-$v_{j}$ $\geq 0\ \forall\,j\in[n]$
-The constraint matrix of $(\mbox{{Asg}}^{\prime})$ can now be used to define a zero-sum game, the hide-and-seek game by von Neumann.
-The game matrix has dimension $2n\times n^{2}$, with only $2n^{2}$ of its $2n^{3}$ entries being nonzero.
-The nonzero entries correspond exactly to the surplus matrix of the assignment problem.
-Recall that we can view the hide-and-seek game as the row and column players picking a vertex and edge of the assignment graph, respectively.
-The column player tries to hide in the edge and the row player tries to search edges by inspecting end points.
-Our correspondence explains why the payoffs of the hide-and-seek game are the inverse of the matching surplus of the respective cell.
-This is due to the rescaling of the linear program to rewrite it in the right form for games.
-Moreover, we can also revisit von Neumann’s correspondence of maximum weight matchings and strategies of the column player.
-His main theorem says that the maximin strategies of the column player (von Neumann calls them optimal strategies) are obtained by permutation matrices corresponding to maximum weight matchings.
-More precisely, if we view a permutation as a bijective function $\sigma\colon[n]\to[n]$, then the maximin strategies are convex combinations of strategies with support determined by permutations $\sigma$ that maximize $\sum_{i=1}^{n}\mu_{i\sigma(i)}$.
-Within the support of the matching specified by a permutation, the probabilities $p$ must be distributed such that $p_{1}\frac{1}{\mu_{1\sigma(1)}}=\dots=p_{n}\frac{1}{\mu_{n\sigma(n)}}$.
-[^8] This proportion is captured by von Neumann in the analysis of a so-called $1$-dimensional game that corresponds to the analysis of game play given a fixed permutation matrix (or, equivalently, perfect matching).
-Hence, it is again the rescaling to obtain $(P^{\prime},D^{\prime})$ that tells how feasible solutions of the primal will be skewed to obtain maximin strategies of the game.
-Once we consider linear programs of the form $(P^{\prime},D^{\prime})$, the correspondence of optimal primal-dual pairs and maximin strategies requires no more skewing but simply multiplication with a scalar.
-Put differently, the scaled matrix leads to a uniform scale for the payoffs.
-5.2 Production Planning Game
-We now provide a more general, natural interpretation of the reduced game for any linear program with positive $b$ and $c$.
-Such linear programs can be viewed as planning optimal production.
-A central authority decides how much of each of $n$ goods to produce, where $x_{j}$ is the amount of good $j$ produced.
+4 Discussion
+The question of generalizing von Neumann’s result (von Neumann, 1953) to an arbitrary primal-dual pair of LPs is clearly an important one.
+Our work explores an extension based on rescaling, the crucial idea that helps define the payoffs in the hide-and-seek game.
+While at first glance this seems only possible if $b>\mathbf{0}$ and $c>\mathbf{0}$, we managed to reduce the case $A\geq\mathbf{0}$ to this case as well.
+Going further may well require a new approach.
+We conclude our paper with discussions tying up a few loose ends.
+4.1 Production Planning Games
+We want to demonstrate the applicability of our reductions by discussing an interpretation of the reduced game for any LP with positive $b$ and $c$ and demonstrate how von Neumann’s hide-and-seek game is a special case.
+LPs with positive $b$ and $c$ can be viewed as planning optimal production: A central authority decides how much of each of $n$ goods to produce, where $x_{j}$ is the amount of good $j$ produced.
 The profit from producing one unit of good $j$ is $c_{j}$.
-The production of each good consumes certain amounts of $m$ resources representing, e.g., raw materials and electricity, or, more broadly, labor hours and machine capacity.
+The production of each good consumes certain amounts of $m$ resources representing, e.g., raw materials and electricity, or even labor hours and machine capacity.
 Producing a unit of the $j$th good requires $a_{ij}$ units of the $i$th resource.
@@
 Selecting the $j$th good and $i$th resource leads to a payment by the column player to the row player of $\frac{a_{ij}}{b_{i}c_{j}}$.
-This quantity is the resource share needed to produce a unit of the $j$th good, $\frac{a_{ij}}{b_{i}}$, divided by the profit $c_{j}$ for producing one unit.
+This quantity is the resource share needed to produce a unit of the $j$th good, that is, $\frac{a_{ij}}{b_{i}}$, divided by the profit $c_{j}$ for producing one unit.
 Together, this is the required resource share per unit profit.
@@
 Hence, it makes intuitive sense that the payoff depends on the reciprocals of both the resource availability and the price of the good.
-5.3 Playing Linear Programs
-We want to broaden the insights of the discussed applications to give an even more general interpretation of our correspondence.
-For comparison, we first recall the original construction by Dantzig (1951) as defined in Section 3 given by the game matrix in Equation 1.
-The skew-symmetric matrix of this game leads to symmetric roles for both players.
-They both specify strategies of the form $(p,q,t)$, where $q$ corresponds to a solution of the primal linear program $(P)$ and $p$ corresponds to a solution of the dual linear program $(D)$.
-By contrast, consider now our construction for positive $b$ and $c$ as detailed in Theorem 4.3.
-There, the column and row players have different roles corresponding to the primal and dual linear programs, respectively.
-Moreover, each fixed strategy gives a guarantee to a player, i.e., a minimum reward for the row player against adversarial play of the column player or a maximum payment by the column player against an adversarial row player.
-These guarantees are obtained by solving $(\mbox{{Row}})$ and $(\mbox{{Col}})$ when fixing strategies $p$ and $q$, respectively.
-When players play optimal strategies, they get their optimal guarantee, which is exactly the maximin or minimax value.
-Importantly, such guarantees are obtained for any strategy, not only when players play optimally and they carry information about the underlying LP.
-First, if the column player can play a (not necessarily optimal) strategy $q$ that guarantees her negative utility, i.e., for this fixed $q$, the minimum $v$ of $(\mbox{{Col}})$ is negative, then this $q$ gives rise to an unboundedness certificate as in Part 1-(b) of the proof of Theorem 4.3.
-Interestingly, the unboundedness can not only be certified by an optimal strategy but by any strategy of the column player yielding nonpositive payoff.
-Conversely, any strategy that guarantees a positive payoff leads to a feasible solution to the corresponding linear program.
-Specifically, assume that $\bar{q}$ and $\bar{p}$ are strategies for the column and row players such that $(\mbox{{Col}})$ and $(\mbox{{Row}})$ have positive optimal values $\bar{v}$ and $\bar{u}$ when fixing $\bar{q}$ and $\bar{p}$, respectively.
-Then, as in Part 1-(b) of the proof of Theorem 4.3, scaling these strategies as $\frac{1}{\bar{v}}\bar{q}$ and $\frac{1}{\bar{u}}\bar{p}$ leads to feasible solutions of $(P^{\prime})$ and $(D^{\prime})$, respectively.
-Their objective values with respect to $(P^{\prime})$ and $(D^{\prime})$ are $\frac{1}{\bar{v}}$ and $\frac{1}{\bar{u}}$.
-Once again, this works even for nonoptimal strategies of the reduced games as long as they guarantee a positive payment.
-However, importantly, such a strategy exists for the row player only if $(D^{\prime})$ admits a feasible solution, and, therefore, $(P^{\prime})$ is not unbounded.
-All these considerations give a broader interpretation of the game play as playing a linear program: In the first case, the column player plays an infeasibility certificate.
-In the second case, the row player can play a feasible strategy, implying that any strategy of the column player has to lead to a positive value of $(\mbox{{Col}})$.
-Hence, in this case, both players play a feasible solution.
-This solution might, however, lead to a suboptimal objective value of the LP.
-In the special case where both play maximin strategies, their play corresponds to playing optimal solutions of the LP, which we can extract alongside the optimal value of the LP.
-Notably, as in the case of the assignment problem, the direct correspondence, where solutions only differ by scaling, is obtained between the linear program pair $(P^{\prime},D^{\prime})$ and the zero-sum game based on the constraint matrix of this game.
-In other words, linear programs of the form $(P^{\prime},D^{\prime})$ are linear programs of games.
-Von Neumann’s result can therefore be interpreted as the realization that the assignment problem can be written in game form.
-And more generally, Proposition 4.1 shows that, whenever $b>\mathbf{0}$ and $c>\mathbf{0}$, any linear program pair of the form $(P,D)$ can be written in game form, as well.
-6 Relationship with Dantzig’s Reduction
-Recall our discussion in Section 1 that Dantzig’s original reduction does not work for the full generality of linear programs.
+This perspective also offers a novel view on the hide-and-seek game.
+In this context, the production planner generates edges.
+For instance, the planner could be an event manager organizing tennis matches.
+A match between players $i$ and $j$ yields a surplus $\phi_{ij}$.
+The resource provider supplies the players, who correspond to the vertices in the assignment problem.
+By the rules of the hide-and-seek game, if a player participates in a desired match, then a compensation of $1/\phi_{ij}$ must be paid.
+Assuming that the surplus from a match accrues linearly over its duration, with a total surplus of $\phi_{ij}$, this compensation corresponds to the fraction of the match that yields one unit of profit.
+4.2 Computational Considerations
+The key idea of the papers repairing Dantzig’s original reduction is to modify his reduced game: Adler (2013) and von Stengel (2023) add an additional row while Brooks and Reny (2023) transform the game with elementary column operations.
+These reductions have in common the fact that the constraint matrix or vectors in the reduced games need large numbers, dependent on the input numbers and exponential in $m$ or nn.
+[^4] As an example, in the reduction by Adler (2013), the additional entry of the constraint vector is estimated as a number that is at least $m(\sqrt{m}\theta)^{m}+1$ where $\theta$ is the absolute value of the largest number contained in $A$ and $b$.
+Nonetheless, since the large numbers can be computed with a polynomial number of arithmetic operations and have a bit size polynomial in the input bit size, the reductions are strongly polynomial.
+[^5] For a formal definition of strongly polynomial time, we refer to Grötschel et al.
+(1993).
+In addition, Adler (2013) also provides a Turing reduction based on sequential variable elimination that avoids large numbers.
+Our reduction is a strongly polynomial reduction without large numbers.
+Within the different realm of computational models based on Turing machines, all of the aforementioned reductions are log-space reductions, i.e., they only use logarithmic space except for outputting the polynomially large reduced instance (Arora and Barak, 2009).
+In fact, the occurring large numbers can be computed bit-wise in logarithmic space.
+This has the interesting consequence that computing maximin strategies in zero-sum games is P-complete because solving linear programs is P-complete (Dobkin et al., 1979).
+Apparently, this observation was missed in previous work.
+4.3 Relationship with Dantzig’s Reduction
+The reduction in the seminal paper by Dantzig (1951) is based on the following skew-symmetric game matrix, given the input data of the primal LP (P)(P).
+[^6] Dantzig (1951) attributes the idea for this matrix to Tucker.
+$\begin{pmatrix}\vskip 6.0pt plus 2.0pt minus 2.0pt\mathbf{0}&A&-b\\ -A^{\intercal}&\mathbf{0}&c\\ b^{\intercal}&-c^{\intercal}&0\\ \end{pmatrix}\text{.}$ (5)
+Denoting a maximin strategy as $(p^{*},q^{*},t^{*})$, the key result of the paper by Dantzig (1951) states:
+  1. 1.
+[…]
+  2. 2.
+    If $b^{\intercal}p^{*}-c^{\intercal}q^{*}<0$, then either $z^{*}=p^{*}$ satisfies (2) or $w^{*}=q^{*}$ satisfies (3) in Proposition 2.1.
+Note that since $(p^{*},q^{*},t^{*})$ is a maximin strategy for a game of value 0, it holds that $t^{*}(b^{\intercal}p^{*}-c^{\intercal}q^{*})=0$.
+The goal of solving $(P)$ is achieved, except for the case $t^{*}=b^{\intercal}p^{*}-c^{\intercal}q^{*}=0$.
+Notably, this “hole” in the reduction was clearly recognized by Dantzig (1951).
 In this section, we explore whether the restrictions that we impose on linear programs suffice to make his reduction work.
 First, we show that under the assumption $b>\mathbf{0}$ and $c>\mathbf{0}$, his reduction works.
-Let $(p^{*},q^{*},t^{*})$ be the minimax strategy of the row player for Dantzig’s game as presented in Equation 1.
+Let $(p^{*},q^{*},t^{*})$ be a minimax strategy of the row player.
 Then, it holds that
-$(i)$ $Aq^{*}-bt^{*}\leq\mathbf{0},$
-$(ii)$ $-A^{\intercal}p^{*}+ct^{*}\leq\mathbf{0},$
-$(iii)$ $b^{\intercal}p^{*}-c^{\intercal}q^{*}\leq\mathbf{0},$
-$(iv)$ $\mathbf{1}^{\intercal}p^{*}+\mathbf{1}^{\intercal}q^{*}+t^{*}=1,$
-$(v)$ $p^{*}\geq\mathbf{0},\;q^{*}\geq\mathbf{0},\;t^{*}\geq 0,$
-$(vi)$ $t^{*}(b^{\intercal}p^{*}-c^{\intercal}q^{*})=0.$
+$\begin{array}[]{lll}(i)\ Aq^{*}-bt^{*}\leq\mathbf{0},&(ii)\ -A^{\intercal}p^{*}+ct^{*}\leq\mathbf{0},&(iii)\ b^{\intercal}p^{*}-c^{\intercal}q^{*}\leq\mathbf{0},\\ (iv)\ \mathbf{1}^{\intercal}p^{*}+\mathbf{1}^{\intercal}q^{*}+t^{*}=1,&(v)\ p^{*}\geq\mathbf{0},\;q^{*}\geq\mathbf{0},\;t^{*}\geq 0,&(vi)\ t^{*}(b^{\intercal}p^{*}-c^{\intercal}q^{*})=0.\end{array}$
 Observe that, as the game is symmetric, $(p^{*},q^{*},t^{*})$ is also a maximin strategy for the column player, and the value of the game is 0.
 Proposition 4.1.
-The following statements are true:
+Assume that $b>\mathbf{0}$ and $c>\mathbf{0}$ and let $(p^{*},q^{*},t^{*})$ be a minimax strategy of the row player in Dantzig’s game given by Equation 5. Then the following statements are true:
   1. 1.
-    If $t^{*}>0$, then $\left(\frac{1}{t^{*}}q^{*},\frac{1}{t^{*}}p^{*}\right)$ is a primal-dual optimal pair for $(P,D)$.
+    If $t^{*}>0$, then $\left(\frac{1}{t^{*}}q^{*},\frac{1}{t^{*}}p^{*}\right)$ is an optimal primal-dual pair for $(P,D)$.
   2. 2.
@@
   1. 1.
-    We first consider the case $t^{*}>0$.
+    We assume $t^{*}>0$.
 Conditions ($i$), (iiii), and ($v$) imply that $\frac{1}{t^{*}}q^{*}$ and $\frac{1}{t^{*}}p^{*}$ are feasible for $(P)$ and $(D)$, respectively.
@@
   2. 2.
-    Now consider the case $t^{*}=0$.
-Assume for contradiction that $q^{*}=0$.
-Then, by ($iv$) and ($v$), it holds that $0\neq p^{*}\geq 0$.
+    Now assume $t^{*}=0$.
+Assume for contradiction that $q^{*}=\mathbf{0}$.
+Then, by ($iv$) and ($v$), it holds that $0\neq p^{*}\geq\mathbf{0}$.
 Hence, since $b>\mathbf{0}$, it follows that $b^{\intercal}p^{*}>0$, contradicting (iiiiii).
 Thus, $q^{*}\neq 0$ which implies (as $c>\mathbf{0}$) that $c^{\intercal}q^{*}>0$.
-Moreover, by ($i$) and ($v$), it holds that $Aq^{*}\leq 0$ and $q^{*}\geq 0$.
+Moreover, by ($i$) and ($v$), it holds that $Aq^{*}\leq\mathbf{0}$ and $q^{*}\geq\mathbf{0}$.
 Together, $q^{*}$ is an unboundedness certificate for (P)(P).∎
 Notably, the second part of Proposition 4.1 has the precondition $t^{*}=0$ whereas the theorem by Dantzig (1951) had the precondition $b^{\intercal}p^{*}-c^{\intercal}q^{*}<0$.
-Recall that the “hole” in Dantzig’s proof was due to the fact that it could happen that both $t^{*}=0$ and $b^{\intercal}p^{*}-c^{\intercal}q^{*}=0$.
+Recall that the issue with Dantzig’s proof was that it could happen that both $t^{*}=0$ and $b^{\intercal}p^{*}-c^{\intercal}q^{*}=0$.
 This raises the question whether Proposition 4.1 works because this case never occurs or whether it works because this case can be dealt with successfully (unlike in Dantzig’s reduction).
@@
 $A=\begin{pmatrix}-1&\frac{1}{2}\\ 1&-\frac{1}{2}\end{pmatrix}\text{, }b=\begin{pmatrix}1\\ 1\end{pmatrix}\text{, and }c=\begin{pmatrix}1\\ 1\end{pmatrix}\text{.}$
-The matrix of Dantzig’s reduced zero-sum game is given by the matrix
-$\begin{pmatrix}0&0&-1&\frac{1}{2}&-1\\ 0&0&1&-\frac{1}{2}&-1\\ 1&-1&0&0&1\\ -\frac{1}{2}&\frac{1}{2}&0&0&1\\ 1&1&-1&-1&0\\ \end{pmatrix}\text{.}$
 […]
@@
 Example 4.3.
-Consider any constraint matrix $A\geq\mathbf{0}$ and consider the linear program pair given by $A$, $b=\mathbf{0}$, and $c=\mathbf{0}$. Dantzig’s reduced game as in Equation 1 for this case is
+Consider any constraint matrix $A\geq\mathbf{0}$ and consider the linear program pair given by $A$ as well as $b=\mathbf{0}$ and $c=\mathbf{0}$.
+Dantzig’s reduced game as in Equation 5 for this case is
 $\begin{pmatrix}\mathbf{0}&A&\mathbf{0}\\ -A^{\intercal}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}^{\intercal}&\mathbf{0}^{\intercal}&0\\ \end{pmatrix}\text{.}$
 […]
-Arguably, the case of $b=c=\mathbf{0}$ in the last example seems somewhat trivial.
+Arguably, the case of $b=\mathbf{0}$ and $c=\mathbf{0}$ in the last example may seem unsatisfying.
 However, we want to stress that Dantzig’s correspondence can only fail for somewhat trivial cases, namely cases covered by our preprocessing in Algorithm 1.
-Indeed, instead of using our reduction, we can also use Dantzig’s reduction in Algorithm 1 for tackling LPs with $A\geq\mathbf{0}$.
+Indeed, instead of using our reduction, we can also use Dantzig’s reduction in algorithm 1 of Algorithm 1 for tackling LPs with $A\geq\mathbf{0}$.
 This approach is sound due to Proposition 4.1.
-7 Computational Considerations
-Up to this point, we have omitted the computational properties of reductions and the number spaces to which they are applicable. In this section, we address these aspects.
-7.1 Reasonable Reductions
-Typically, in complexity theory, meaningful reductions are required to be polynomial, that is, the effort of the reduction, as well as the effort to recover the solution, is required to be polynomial in the size of the problem.
-We shall assume for now that the data of the linear programs is rational and discuss implications of our reduction for real input data subsequently.
-To make the reduction meaningful in the framework of complexity theory and given that any LP problem with rational data is known to be solvable in time polynomial in the size of the problem, we follow Adler (2013) and require reductions to be strongly polynomial, that is, they have to use a polynomial number of arithmetic operations and, if applied to rational input data, polynomial space (Tardos, 1986; Grötschel et al., 1993).
-In our case, the input data of a linear program depends on its dimensions $m$ and $n$ and the encoding of numbers, hence polynomial means with respect to $m$, $n$, and the logarithm of numbers such as the entries of the constraint matrix.
-Note that strongly polynomial reductions still permit the occurrence of exponentially large numbers with respect to $m$ and $n$, and hence all the previously existing reductions by Adler (2013), Brooks and Reny (2023), and von Stengel (2023) are strongly polynomial as well.
-In the context of linear programming, strongly polynomial reductions make a lot of sense since it is a long-standing open problem to decide whether linear programs can be solved in strongly polynomial time (see, e.g., Tardos, 1986; Grötschel et al., 1993, Section 6.6).
-Notably, our as well as previous reductions from linear programs to zero-sum games are even log-space reductions, i.e., they only use logarithmic space except for outputting the polynomially large reduced instance (Arora and Barak, 2009).
-Note that even a log-space reduction can produce exponentially large numbers with respect to input parameters, such as the large numbers in the reductions by Adler (2013), Brooks and Reny (2023), and von Stengel (2023).
-Their reductions run in log-space because these numbers do not have to be computed with arithmetic operations but can directly be inferred from the input parameters and outputted without computation.
-[^9] For example, recall the number $m(\sqrt{m}\theta)^{m}+1$ from Footnote 4.
-In binary representation, this number only uses a polynomial number of bits with respect to the input size.
-Instead of being computed it can be inferred directly from the input data and simply outputted as a matrix entry of the reduced game.
-An interesting feature of previous reductions being log-space is that it implies that computing maximin strategies in zero-sum games is P-complete because solving linear programs is P-complete (Dobkin et al., 1979).
-Apparently, this observation was missed in previous work.
-7.2 Reductions for Real Numbers
-In the previous subsection, we assumed rational data for the discussion of reasonable reductions.
-However, in the beginning of Section 2, the problems are presented in terms of real numbers.
-This presentation is common whenever the mathematics of linear programming problems are discussed.
-The reason for this is that, as recognized from the early days of linear programming and game theory, all major mathematical theorems (e.g., the LP-Duality Theorem or the Minimax Theorem) are valid for real numbers.
-[^10] In fact, they are valid for any ordered field (see, e.g., Gale, 1960)).
-In addition, most related algorithms, such as Fourier-Motzkin elimination or the simplex algorithm, prior to the polynomial-time algorithm for solving LPs by Khachiyan (1979) were developed and presented in terms of real numbers.
-Indeed, both the reductions by Dantzig (1951) and von Neumann (1953) only use elementary computations and are applicable for real numbers.
-However, the corrected extensions of Dantzig’s reduction of Adler (2013), Brooks and Reny (2023), and von Stengel (2023) are only applicable for rational LPs because the constructions require a big number depending on the data.
-It is currently not known whether this number can be calculated in strongly polynomial time for real-valued LPs.
-[^11] By contrast, the Turing-reduction by Adler (2013) is applicable for real input but it requires solving up to $n$ LPs.
-Not being applicable to real numbers feels like a missing feature of these reductions.
-LPs and zero-sum games in real numbers are generally relevant for complexity analysis of computational models with real numbers (see, e.g., Blum et al., 1989).
-In particular, the problem of finding a strongly polynomial algorithm for real-valued LPs is listed as one of the most prominent mathematical challenges for the 21st century (Smale, 1998, Problem 9).
-Interestingly, our reduction shares the applicability for real numbers with the original reduction of von Neumann.
-It thus provides a somewhat more complete equivalence between LPs and zero-sum games, albeit only for positive $b$ and $c$ or nonnegative $A$.
-Note that only reductions from linear programs to zero-sum games that are applicable to real numbers can be used to translate strongly polynomial algorithms (for real input) for finding maximin strategies for zero-sum games to strongly polynomial algorithms to solve linear programs.
-In other words, using our technique, any strongly polynomial algorithm for computing maximin strategies for zero-sum games would immediately give rise to a strongly polynomial algorithm for solving linear programs with positive $b$ and $c$ or nonnegative $A$.
-8 Discussion
-Clearly, the question of generalizing von Neumann’s result (von Neumann, 1953) to an arbitrary primal-dual pair of LPs is an important one.
-Our work explores an extension based on rescaling, the crucial idea that helps define the payoffs in the hide-and-seek game.
-Even though this seems only possible if $b>\mathbf{0}$ and $c>\mathbf{0}$, we managed to reduce the case $A\geq\mathbf{0}$ to this case as well.
-Going further may well require a new approach.
-Our paper is first and foremost an attempt at expounding a purely technical result; yet, its historical overtones are hard to miss.
-In that vein, we would like to end with some observations which may interest students and observers of the history of computing and computability.
-A lot has been said about von Neumann’s extraordinary intellectual prowess and his foundational influence on the creation of the digital computer, in particular, his 1945 report on what is known today as the von Neumann architecture (see, e.g., Halmos, 1973; Bhattacharya, 2021).
-Our study of his masterful work (von Neumann, 1953) has led us to believe that he had an intuitive understanding of the notion of efficient computability, i.e., polynomial-time solvability.
-With a view to solving the assignment problem efficiently, von Neumann (1953) gave the following procedure: Reduce the assignment problem to a zero-sum game and approximately solve the latter using methods from Brown and von Neumann (1951), involving solving a sequence of differential equations.
-Over the centuries, other mathematicians had also found algorithms that were later seen to be polynomial-time algorithms, e.g., Euclid’s algorithm for computing the greatest common divisor of two numbers.
-However, von Neumann’s work is qualitatively different.
-He explicitly reasoned about the time complexity of the method he had proposed, as is evident from reading the first page of the paper (von Neumann, 1953).
-The page starts by observing that an optimal solution to the assignment problem can be found via an exhaustive search, i.e., by considering all $n!$ permutation matrices, and ends by stating that his procedure gives a “significant approximate solution” in “a moderate power of nn” steps.
-The page ends with the following striking sentence, which is clearly alluding to the qualitative difference between polynomial and exponential time.
-It turns out that this number is a moderate power of $n$, i.e., considerably smaller than the “obvious” estimate $n!$ mentioned earlier.
-Additionally:
-  1. 1.
-    Von Neumann (1953) reduced the assignment problem to the problem of finding a Nash equilibrium in zero-sum games.
-  2. 2.
-    Today we know that both these problems are in P: the first due to the work of Kuhn (1955) and the second because the problem is equivalent to linear programming, which is in P due to Khachiyan (1979); moreover, the reduction is polynomial time.
-Recall that the class P was defined more than a decade after von Neumann’s work (Edmonds, 1965).
-Hence, von Neumann gave a strongly polynomial-time reduction between two problems in P — over a decade before P was even defined!
 Acknowledgements
+We would like to thank the editor and anonymous referees from GEB for their excellent comments.
 Most of this research was done while Martin Bullinger was affiliated with the University of Oxford.